### v1

# Reduced Order Hysteretic Magnetization Model for Composite Superconductors

Abstract
In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors. Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without having to solve for the detailed current density distribution. It can be used as a homogenized model in large-scale superconducting systems in order to significantly reduce the computational work compared to detailed simulations. In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to an external transverse magnetic field. We propose two levels of ROHM models: a rate-independent model that reproduces pure hysteresis without dynamic effects, and a rate-dependent model that generalizes the former by also reproducing dynamic effects observed in superconducting strands due to coupling and eddy currents. We describe the implementation and inclusion of these models in a finite element framework, discuss their parameter identification and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency.
  • September 20, 2024
Keywords: Reduced Order Method, Hysteresis Model, AC Loss, Magnetization.
[…]
*                                    This work is a preliminary version. Copyright may be transferred without notice, after which this version may no longer be accessible.
[…]
*                                    This work is a preliminary version. Copyright may be transferred without notice, after which this version may no longer be accessible.
1 Introduction
Large-scale superconducting systems are multi-scale structures made of a high number of wounded superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes. The magnetic response of these systems is hysteretic and rate-dependent. Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting parts [1, 2, 3, 4], whereas rate-dependency is caused by eddy currents, and coupling current dynamics between different conductors [5, 6, 7].
Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8]. Their computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
Most approaches to describe the superconducting hysteresis are based on the definition of a relationship between electric field and current density, either non-smooth, as in the Bean model [1], or smooth, as in the power law model [9]. The use of the power law model leads to very accurate results and versatile numerical models with, e.g., the finite element (FE) method [10], which can also be combined with eddy current and coupling current models.
Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes the numerical models computationally demanding to solve, already for single strands or cables [11]. As a consequence, modelling large-scale superconducting systems in all their details down to the conductor level is completely unrealistic in practice. Faster, more efficient, methods are necessary.
Homogenization methods which reproduce the macroscopic effects of coupling dynamics without modelling them explicitly are good candidates. The general idea of these methods is to describe the systems in terms of the average fields [12], in order to reproduce the magnetization and loss accurately, without having to solve for the detailed current density distribution, which has the potential to strongly reduce the computational cost of the simulations.
Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures [13, 14, 15] with bundles of wires of any shape, described in the frequency or time domain [16, 17]. They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions, or on experimental measurement results.
Following a similar approach for superconducting systems requires to use dedicated history-dependent parameters in order to reproduce their hysteretic response.
A few hysteresis models have been proposed for modelling superconductor. A vector model is proposed in [8, 18] for field quality computation in accelerator magnets. This model describes the magnetization of superconducting filaments using nested magnetization ellipses. Another hysteresis model based on the Preisach model [19] was proposed in [20, 21] in the form of an equivalent circuit element. Yet another vector model based on a variational approach was described in [22]. These models however only describe rate-independent hysteresis and are therefore not suited for modelling transient effects in composite superconductors.
The literature on hysteresis models for ferromagnetic materials is much more extensive. Among the most popular models are the Jiles-Atherton [23], Preisach [19] models, and rate-dependent models such as the Chua-Stromsmoe model [24, 25]. In 1997, Bergqvist proposed a thermodynamically consistent approach to model hysteresis [26], which led to the development of the energy-based model [27, 28, 29].
The energy-based model offers several advantages. First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions. Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that can be combined. Third, it is consistently defined as a vector, and not scalar, model from the beginning [27]. Fourth, it can be directly generalized to include rate-dependent effects [28]. Finally, the equations of this model consist in simple explicit tests, which lead to straightforward implementations and very efficient simulations.
In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis by focussing on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis. This situation is relevant for most applications and constitutes a first step towards the homogenization of a complete magnet winding. This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects: eddy currents in normal conductors and interfilamentary coupling currents. For this paper, we assume that the strand carries no transport current. The inclusion of transport current will be the focus of further work. We also consider a constant temperature and do not model thermal effects.
We propose two models: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the first model by also describing the eddy current and coupling current effects, in addition to the superconducting hysteresis.
The models are general and can be adapted to a variety of composite superconductors, such as strands, tapes, or cables. They can also be easily implemented in any FE framework allowing for user-defined material properties.
This paper is structured as follows. In Section 2, we introduce the model of a multifilamentary strand and describe its response in terms of magnetization and loss for transverse fields in a range of frequencies and amplitudes. In Section 3, we present the equations leading to rate-independent and rate-dependent hysteresis models. In Section 4, we discuss the implementation of the equations and describe how to include them in a two-dimensional (2D) FE framework with a $\phi$-formulation. We then propose in Section 5 a parameter identification procedure for the models. Finally, in Sections 6 and 7, we apply the hysteresis models in rate-independent and rate-dependent situations, respectively, and demonstrate their potential in the context of homogenization techniques.
2 Multifilamentary Strand Dynamics
In this section, we present the response of a composite multifilamentary superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization. To this end, we model the detailed current density distribution inside the strand with a FE method and we analyze the different dynamics that result from the strand composite structure.
2.1 Problem definition
We consider a strand of diameter $d$ represented in Fig. 1. It consists of $N_{\text{f}}$ Nb-Ti filaments twisted with twist pitch length $p$ and embedded in a conducting copper (Cu) matrix. The geometrical parameters of the strand are summarized in Table 1. The strand carries no transport current but is subject to an applied (app) transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of amplitude $h_{\text{max}}$ (A/m), frequency $f$ (Hz), and direction $\hat{\boldsymbol{e}}_{y}$.
Figure 1: Composite strand with Nb-Ti filaments embedded and twisted in a copper matrix (cylindrical outline) subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)$.
Number of filaments ($N_{\text{f}}$) 54
Filament diameter $90~{}\upmu$m
Filament center-to-center spacing $110~{}\upmu$m
Strand diameter ($d$) $1$ mm
Twist pitch length ($p$) $19$ mm
Table 1: Geometrical parameters of the strand.
We fix the temperature to $1.9$ K. The resistivity of the Nb-Ti filaments is described by the power law [9] with power index $n=30$ and the field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$ (A/m2) given in Fig. 2, with $\boldsymbol{b}$ (T) the local magnetic flux density, defined as $\boldsymbol{b}=\mu_{0}\boldsymbol{h}$, with $\mu_{0}=4\pi\times 10^{-7}$ H/m the permeability of vacuum. The resistivity of the copper matrix, accounting for magneto-resistance, is obtained from the STEAM material library [30], with a residual resistivity ratio $\text{RRR}=100$.
Figure 2: Field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$. The maximum value is $j_{\text{c}}(\boldsymbol{0})=2.4\times 10^{10}$ A/m2.
We model the strand response with the CATI method, proposed in [31]. This method accounts for 3D effects due to the strand twist by using a pair of 2D models solved on a cross section of the strand, and coupled via circuit equations. As it is based on 2D models, it offers fast simulations. The method is implemented in GetDP [32] within FiQuS [33]. FiQuS is developed at CERN as part of the STEAM framework [34]. Geometry and mesh are performed by Gmsh [35]. All the software is open-source and free to use. The code is available online [^2] https://gitlab.cern.ch/steam/analyses/cati-strand.
2.2 Magnetization and loss
We focus on magnetization and power loss. These are the two quantities we want to reproduce directly with the hysteresis model without having to compute the detailed current density distribution.
The average magnetization vector $\boldsymbol{m}$ (A/m) is defined as the magnetic dipole moment per unit volume [36]. Based on the current density distribution on a given cross section, we have
$\boldsymbol{m}=2\cdot\frac{1}{2a}\int_{S}\boldsymbol{x}\times\boldsymbol{j}\,\text{d}S,$ (1)
with $a=\pi d^{2}/4$ the surface area of the strand cross section $S$, $\boldsymbol{x}$ the position vector, and $\boldsymbol{j}$ the current density. The multiplication by $2$ accounts for current loops closing at infinity in this 2D problem, as justified in A.
The instantaneous power loss per unit length $P_{\text{tot}}$ (in W/m) is obtained by integrating the instantaneous power loss density, or Joule loss, over the strand cross section, including the Nb-Ti filaments and the Cu matrix,
$P_{\text{tot}}=\int_{S}\boldsymbol{j}\cdot\boldsymbol{e}\,\text{d}S,$ (2)
with $\boldsymbol{e}=\rho\boldsymbol{j}$ the electric field (and $\rho$ the resistivity). The power loss per cycle and per unit length (in J/m) is evaluated as
[…]
To make the interpretation of the loss easier, and also because it will be helpful for the construction of the hysteresis model, the total power loss per cycle $Q_{\text{tot}}$ is decomposed in three contributions, as was done in [31]: hysteresis, coupling, and eddy loss. The hysteresis loss is the loss induced by currents flowing in the superconducting filaments. The coupling loss is the loss induced by currents flowing in the conducting matrix, perpendicular to $\hat{\boldsymbol{e}}_{z}$ (the coupling currents). Finally, the eddy loss is the loss induced by currents flowing in the conducting matrix along $\hat{\boldsymbol{e}}_{z}$.
Magnetization and loss are closely related to each other. Indeed, the area inside a closed magnetization loop is proportional to the total loss per cycle [37, 38]. We have, in terms of the applied magnetic field $\boldsymbol{h}_{\text{app}}$,
$Q_{\text{tot}}=a\oint\mu_{0}\boldsymbol{h}_{\text{app}}\cdot\text{d}\boldsymbol{m}.$ (4)
2.3 Dynamic response of the strand
The strand response is computed by the CATI method for a range of field amplitudes $h_{\text{max}}$ and frequencies $f$. The power loss per cycle induced by the transverse field is represented in Fig. 3 as a function of $f$ for three different values of $h_{\text{max}}$. The $y$-component of the magnetization vector, $m$, is represented in Figs. 4 and 5 for $\mu_{0}h_{\text{max}}=2$ T as a function of $\mu_{0}h_{\text{app}}(t)$ for selected frequencies.
Figure 3: Total loss per cycle and its contributions as a function of the frequency $f$, for three different magnetic field amplitudes. The legend is the same for the three subfigures. Results from [31].
At low frequencies, when the rate of change of the transverse magnetic field is not high enough for significant coupling currents and eddy currents to take place in the matrix, the superconducting filaments behave close to uncoupled. In that regime, most loss is due to hysteresis in the filaments. A typical magnetization loop in that case is represented in Fig. 4. The magnetization decreases when the field increases as a result of the field-dependent critical current density. In this regime, a good first approximation is to consider that the response is rate-independent (it is not strictly the case in reality due to the finite value of the $n$ index in the power law). To this end, a rate-independent hysteresis model can be used.
Figure 4: Strand magnetization along $\hat{\boldsymbol{e}}_{y}$ for $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. Same result as the first curve in Fig. 5.
When the frequency increases, the response of the strand becomes strongly rate-dependent. As shown in Fig. 3, coupling and eddy loss can dominate the total loss, and the filament loss is not constant with frequency. As a result, the loss per cycle can exhibit more than one order of magnitude of variation as a function of frequency for an identical applied field amplitude. Similarly, the magnetization strongly depends on the frequency, as shown in Fig. 5. To reproduce these regimes, there is a clear need for a rate-dependent hysteresis model. The design of such a model benefits from a good understanding of the physics of the curves in Figs. 3 and 5.
Figure 5: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T and for different frequencies $f$. Results from [31] using Eq. (1).
For low field amplitudes, the dynamics of the coupling currents and of the associated loss is well described by analytical models [5, 39]. The coupling loss follows follows a typical bell curve with a maximum at a characteristic frequency $f_{\text{c}}$.
With increasing field amplitude, the peak of the coupling loss curve shifts towards lower frequencies, as can be seen in Fig. 3 for $\mu_{0}h_{\text{max}}=2$ T. This is associated with the saturation of superconducting filaments, which limit the coupling currents amplitude, and to a change of the effective permeability, due to the loss of diamagnetic effect due to saturated filaments, as discussed in [5].
The coupling currents also give rise to a change of regime for the filament magnetization and loss. As mentioned previously, at low frequencies, $f\lesssim f_{\text{c}}$, they are mostly uncoupled, producing a relatively small magnetization. On the other hand, for $f\gtrsim f_{\text{c}}$, they are coupled and exchange currents via coupling currents. As a result, their magnetization is increased, see for example the situation for $f=10$ Hz in Fig. 5. Whether the filament loss for coupled filaments is higher or lower than that for uncoupled filaments depends on $\mu_{0}h_{\text{max}}$, as can be seen in Fig. 3.
At even higher frequencies, eddy currents in the conducting matrix become the dominating factor for loss and magnetization. As a result of eddy currents and the skin effect, the inner part of the strand, containing the filaments, is shielded from the outer field and the filament loss decreases accordingly. In that regime, the magnetization curve approaches the shape of an ellipse [37] that gets thinner and thinner with increasing frequency, and of slope approaching $-2$ because of demagnetization effects (this is further explained in Section 3.1[38].
3 Reduced Order Hysteretic Magnetization Model
The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method. This is fast and convenient enough to get a good understanding of a single strand response. However, for simulating a full-scale magnet, containing thousands of strands, one cannot afford such a detailed description of the current density at the strand level. Instead, one has to use an alternative model that produces similar macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but without requiring a detailed calculation of the current density, so as to strongly reduce the computational cost. This idea is illustrated in Fig. 6. We present such a model in this section, the Reduced Order Hysteretic Magnetization (ROHM) model.
[…]
This model is adapted and inspired from the energy-based model for ferromagnetic materials [27, 28].
We start in Section 3.1 by clarifying how the ROHM model can be used to model the strand. We then define the building blocks, referred to as cells, that will be combined for the construction of the model. Each cell defines a local hysteretic relationship between the magnetic field $\boldsymbol{h}$ and the magnetic flux density $\boldsymbol{b}$. We present in Section 3.2 a rate-independent cell, that we call the superconductor cell (S cell). In Section 3.3, we add contributions to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell). In Section 3.4, we discuss how to reproduce the effect of a field-dependent critical current density within the cells. Finally, we show in Section 3.5 how several cells can be combined into a chains of cell, defining a complete ROHM model.
3.1 Concept of a reduced order model
[…]
$\boldsymbol{b}=\mu_{0}\boldsymbol{h}+\mu_{0}\boldsymbol{m}.$ (5)
Similarly, the power loss is no longer obtained via a current density as in Eq. (2), but is now instead described by the magnetic power density $\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$. Part of this power is related to stored magnetic energy, and hence not associated with loss, and the remaining part is related to dissipated energy. One major advantage of the energy-based hysteresis model used in this work is that it gives access to both parts separately at any time [27].
[…]
$\boldsymbol{h}=\boldsymbol{h}_{\text{app}}-\frac{1}{2}\boldsymbol{m}.$ (6)
[…]
$\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\frac{1}{2}\boldsymbol{m},$ (7)
where $\boldsymbol{h}_{\text{in}}$ has no immediate meaning in the detailed strand model per se, but is the local magnetic field the hysteresis model will see inside the strand. These curves are shown in Fig. 7. At high frequencies ($f\gtrsim 10^{4}$ Hz), they approach the shape of an ellipse with a slope of $-1$. The area of closed loops are identical in both Figs. 5 and 7, as shown in B.
Figure 7: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T. Same results as in Fig. 5 but in terms of $\mu_{0}h_{\text{in}}$ instead of $\mu_{0}h_{\text{app}}$ (see Eq. (7)).
3.2 Superconductor cell - S cell
As a first step, we present a rate-independent hysteresis cell [27], that we refer to as the superconductor cell, or S cell. In this cell, the magnetic field $\boldsymbol{h}$ is decomposed into a reversible field, $\boldsymbol{h}_{\text{rev}}$, and an irreversible field, $\boldsymbol{h}_{\text{irr}}$:
$\boldsymbol{h}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{irr}}.$ (8)
The reversible field defines the magnetic flux density:
$\boldsymbol{b}=\mu_{0}\boldsymbol{h}_{\text{rev}}.$ (9)
The irreversible field creates hysteresis by introducing history dependence. Its amplitude $\|\boldsymbol{h}_{\text{irr}}\|$ is bounded by a value $\kappa$ (A/m), the irreversibility parameter. From a known magnetic field $\boldsymbol{h}$, which is the driving vector, and a given reversible field $\boldsymbol{h}_{\text{rev}}$ that has been established previously (as a function of the history of $\boldsymbol{h}$), the irreversible field $\boldsymbol{h}_{\text{irr}}$ is determined as follows:
$\boldsymbol{h}_{\text{irr}}$ =$&\boldsymbol{h}-\boldsymbol{h}_{\text{rev}},\quad&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|<\kappa,\\ &\kappa\ \dot{\boldsymbol{b}}/\|\dot{\boldsymbol{b}}\|,\quad&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|=\kappa,\\ \end{aligned}\right.$ (10)
where the dot notation (as in $\dot{\boldsymbol{b}}$) represents the time derivative of the quantity. An illustration of this equation is given in Fig. 8. While the driving field $\boldsymbol{h}$ evolves inside the sphere of radius $\kappa$ centered at $\boldsymbol{h}_{\text{rev}}$, the reversible field $\boldsymbol{h}_{\text{rev}}$ stays constant, and the irreversible field is defined accordingly, see Fig. 8(a). If $\boldsymbol{h}$ reaches the surface of the sphere, then, the reversible field $\boldsymbol{h}_{\text{rev}}$ must evolve as well, in order to maintain the condition $\|\boldsymbol{h}_{\text{irr}}\|\leq\kappa$ valid, and $\boldsymbol{h}_{\text{irr}}$ is established parallel to $\dot{\boldsymbol{b}}$, see Fig. 8(b).
(a) $\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|<\kappa$.
(b) $\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|=\kappa$.
Figure 8: S cell, adapted from [27, 29]. Illustration of the magnetic field decomposition into reversible and irreversible fields in 2D. The dotted elements in (b) refer to the situation at a previous time.
The total power density $p_{\text{tot}}$ (in W/m3) is expressed as
$p_{\text{tot}}$ $=\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$
$=\underbrace{\boldsymbol{h}_{\text{rev}}\cdot\dot{\boldsymbol{b}}}_{p_{\text{rev}}}+\underbrace{\boldsymbol{h}_{\text{irr}}\cdot\dot{\boldsymbol{b}}}_{p_{\text{irr}}}.$ (11)
The reversible power $p_{\text{rev}}$ is the time derivative of a stored magnetic energy density,
$p_{\text{rev}}=\boldsymbol{h}_{\text{rev}}\cdot\mu_{0}\dot{\boldsymbol{h}}_{\text{rev}}=\frac{d}{dt}\left(\frac{1}{2}\boldsymbol{h}_{\text{rev}}\cdot\mu_{0}\boldsymbol{h}_{\text{rev}}\right),$ (12)
while the irreversible power $p_{\text{irr}}$ is always non-negative and is associated with dissipated energy density,
$p_{\text{irr}}=\boldsymbol{h}_{\text{irr}}\cdot\dot{\boldsymbol{b}}=\kappa\|\dot{\boldsymbol{b}}\|=\kappa\|\mu_{0}\dot{\boldsymbol{h}}_{\text{rev}}\|.$ (13)
It is one of the main benefits of the energy-based approach that the dissipated energy is clearly separated from the stored energy [27].
A mechanical analogy of a S cell in a one-dimensional setting is shown in Fig. 9. It represents a mechanical system made of a dry friction element in parallel with a restoring spring [27]. Examples of hysteresis curves computed with the S cell are shown in Fig. 10. These curves are independent of the rate of change of the applied field.
Figure 9: Mechanical analogy [27] of the S cell: a parallel connection of a dry friction element and a restoring spring. The system is subject to a force $h$, which causes an elongation $b=\mu_{0}h_{\text{rev}}$.
Figure 10: Hysteresis curves for the S cell subject to a unidirectional oscillating magnetic field $h$ as an input. The legend is the same for both subfigures. (Left) Magnetic flux density $b$ versus applied field $\mu_{0}h$. (Right) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
3.3 Composite superconductor cell - CS cell
In regimes in which the superconducting strand response is clearly not rate-independent, the S cell presented in the previous section must be generalized. Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq. (8) is rewritten
$\boldsymbol{h}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{irr}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}}.$ (14)
The equation for $\boldsymbol{h}_{\text{irr}}$ is modified accordingly:
$\boldsymbol{h}_{\text{irr}}$ =$&\boldsymbol{h}-\boldsymbol{g},\quad&&\text{if }\|\boldsymbol{h}-\boldsymbol{g}\|<\kappa,\\ &\kappa\ \dot{\boldsymbol{b}}/\|\dot{\boldsymbol{b}}\|,\quad&&\text{if }\|\boldsymbol{h}-\boldsymbol{g}\|=\kappa,\\ \end{aligned}\right.$ (15)
with $\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}}$. The magnetic field decomposition is represented in Fig. 11(b). As described below, the fields $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$ are always parallel to $\dot{\boldsymbol{b}}=\mu_{0}\dot{\boldsymbol{h}}_{\text{rev}}$. We refer to this new cell as the composite superconductor cell, or CS cell.
Adapting [28], the field $\boldsymbol{h}_{\text{eddy}}$ is defined as
$\boldsymbol{h}_{\text{eddy}}=\tau_{\text{e}}\dot{\boldsymbol{h}}_{\text{rev}},$ (16)
with an eddy current time constant parameter $\tau_{\text{e}}$ (s). This component produces a loss contribution, $p_{\text{eddy}}$, which is non-negative. It reads:
$p_{\text{eddy}}=\boldsymbol{h}_{\text{eddy}}\cdot\dot{\boldsymbol{b}}=\frac{\mu_{0}}{\tau_{\text{e}}}\|\boldsymbol{h}_{\text{eddy}}\|^{2}=\frac{\tau_{\text{e}}}{\mu_{0}}\|\dot{\boldsymbol{b}}\|^{2}.$ (17)
In the mechanical analogy, the inclusion of eddy current effects corresponds to adding a dashpot (viscous friction element) in parallel with the dry friction element and the restoring spring, as illustrated in Fig. 11(a). For slow variations of the force $h$, the dashpot is barely opposing the elongation $b$, whereas for faster variations of $h$, it becomes gradually more blocking and starts to oppose to variations of $b$. This reproduces magnetic shielding and increased magnetization due to eddy currents.
The next new contribution $\boldsymbol{h}_{\text{coupling}}$ is associated with coupling currents [5]. The coupling currents in the conducting matrix are fed by currents flowing in the superconducting filaments, and are therefore also subject to hysteresis and saturation of the filaments. The field $\boldsymbol{h}_{\text{coupling}}$ has then its own irreversibility parameter $\chi$ (in A/m). We define
$\boldsymbol{h}_{\text{coupling}}$ =$&\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}},&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|<\chi,\\ &\chi\ \dot{\boldsymbol{h}}_{\text{rev}}/\|\dot{\boldsymbol{h}}_{\text{rev}}\|,&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\geq\chi.\\ \end{aligned}\right.$ (18)
$=\frac{\tau_{\text{c}}}{\mu_{0}}\dot{\boldsymbol{b}}_{\text{coupling}},$ (19)
with a coupling currents time constant $\tau_{\text{c}}$ (s). The field $\boldsymbol{h}_{\text{coupling}}$ introduces two new loss contributions, $p_{\text{coupling}}$ and $p_{\text{irr,c}}$, that read:
$p_{\text{coupling}}$ $=\frac{\mu_{0}}{\tau_{\text{c}}}\|\boldsymbol{h}_{\text{coupling}}\|^{2}=\frac{\tau_{\text{c}}}{\mu_{0}}\|\dot{\boldsymbol{b}}_{\text{coupling}}\|^{2},$ (20)
$p_{\text{irr,c}}$ $=\boldsymbol{h}_{\text{coupling}}\cdot\left(\dot{\boldsymbol{b}}-\dot{\boldsymbol{b}}_{\text{coupling}}\right).$ (21)
The mechanical analogy associated with this new problem is represented in Fig. 11(a). The combination of the dashpot in series with the dry friction element reproduces the saturation of the coupling currents. When the force on the dashpot remains below $\chi$, the dry friction element is at rest and produces no hysteresis. This corresponds to a situation in which the coupling currents are small enough for the hysteresis to be neglected. If the force exceeds $\chi$ on the dashpot, it is then subject to hysteresis. The second hysteresis element, with $\chi$, always acts in parallel with the first one, related to filament magnetization. The magnetization of coupled filaments therefore depends of the sum of both.
(a) Mechanical analogy.
(b) Illustration in 2D.
Figure 11: CS cell accounting for eddy currents and coupling current effects. The dashed circle of radius $\chi$ limits the amplitude of the field $\boldsymbol{h}_{\text{coupling}}$.
Note that by choosing $\tau_{\text{e}}=0$ and $\tau_{\text{c}}=0$, we retrieve the S cell of Section 3.2. The S cell is therefore a particular case of the CS cell.
An illustration of magnetization curves obtained with a CS cell is given in Fig. 12, and the loss per cycle as a function of frequency is given in Fig. 13. The loss per cycle curve exhibits the sought features for the composite strand. The filament curve shows the expected evolution, two plateaus at low and middle frequencies, with a transition where the coupling loss peaks, and then a sharp decrease when the eddy loss becomes dominant. Coupling and eddy loss curves both follow the shape of two bell curves.
Figure 12: Hysteresis curves for the composite hysteresis cell reproducing eddy and coupling currents, with $\kappa=0.2/\mu_{0}$ A/m, $\chi=0.3/\mu_{0}$ A/m, $\tau_{\text{c}}=0.1$ s, and $\tau_{\text{e}}=70$ $\upmu$s. Unidirectional oscillating magnetic field $h$ as an input. The legend is shared between both subfigures. (Left) Magnetic flux density $b$ versus applied field $\mu_{0}h$. (Right) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
Figure 13: Total loss per cycle and separated loss contributions as a function of the frequency $f$, for a magnetic field amplitude of $2$ T obtained with the same composite hysteresis cell as in Fig. 12.
Note that the eddy current curve for the eddy loss per cycle should be asymptotically decreasing as $1/\sqrt{f}$ at high frequencies, due to the skin effect. With the CS cell, one can however only expect the curve to decrease as $1/f$ at high frequencies. For $f\to\infty$, the CS cell is therefore not expected to provide perfectly valid results. It should be further improved if a more realistic frequency response is wanted in such high frequencies. For most practical applications however, the field rates of change are sufficiently low for this error to stay small.
Comparing with the results from a detailed multifilamentary strand model in Fig. 3, one can see that a single CS cell is still not sufficiently accurate. Several of them have to be combined to achieve a better matching with the reference, as will be discussed in Section 3.5.
3.4 Inclusion of $j_{\text{c}}(b)$ field dependence
In reality, the magnetization of superconductors decreases with increasing field, as a consequence of the decreasing critical current density $j_{\text{c}}$. Such an effect is not yet reproduced with the hysteresis cells defined above, because the irreversibility parameters $\kappa$ and $\chi$ are still constant values. By making the irreversibility parameters $\kappa$ and $\chi$ decrease with increasing field, one can produce more realistic magnetization curves.
In general, we therefore define field-dependent irreversibility parameters:
$\kappa(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa},\quad\chi(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi},$ (22)
with $\bar{\kappa}$ and $\bar{\chi}$ two constants (in A/m) and $f_{\kappa}$ and $f_{\chi}$ two scaling functions, smoothly evolving from $1$ at zero field to 00 at large fields. The expressions of $f_{\kappa}$ and $f_{\chi}$ are functions of the geometry and must be determined at the parameter identification step. This will be discussed in Section 5.
Note that the introduction of field-dependent irreversibility parameters in the cells brings an additional nonlinearity to the equations to be addressed during the numerical simulation.
3.5 Chain of cells
The S cell and the CS cell contain the necessary ingredients to reproduce the different regimes observed in the multifilamentary strand, but a single cell alone does not provides a faithful description over a wide field range. Indeed, if the applied magnetic field is smaller than $\kappa$, both the S cell and the CS cell produce exactly zero magnetic flux density, and hence no loss. In reality, the evolution towards saturation is smooth and progressive, and not subject to a single threshold $\kappa$, as in Figs. 10 and 12.
To better approach the smooth magnetization curves of Figs. 4 and 5, we can combine several cells into what we call a chain of cells. To this end, we follow the approach proposed in [27].
The approach consists of decomposing the total magnetic flux density $\boldsymbol{b}$ into a number of $N\in\mathbb{N}_{0}$ contributions $\alpha_{k}\boldsymbol{b}_{k}$, with $k=1,\dots,N$. The $\boldsymbol{b}_{k}$ are described by distinct hysteresis cells, with distinct parameter values and the $\alpha_{k}$ are weights associated with each of them. All cells are subject to the same total magnetic field $\boldsymbol{h}$ and contribute to the total magnetic flux density with weight $\alpha_{k}$. We define
$\boldsymbol{b}=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{b}_{k}=\sum_{k=1}^{N}\alpha_{k}\mu_{0}\boldsymbol{h}_{\text{rev},k},\ \quad\sum_{k=1}^{N}\alpha_{k}=1,$ (23)
where $\boldsymbol{h}_{\text{rev},k}$ is the reversible field associated with cell $k$. With the mechanical analogy, this corresponds to connecting cells as a chain, i.e., in series, as illustrated in Fig. 14. The power loss is computed cell by cell, in terms of the fractions $\alpha_{k}\dot{\boldsymbol{b}}_{k}$ of the total rate of change $\dot{\boldsymbol{b}}$.
Figure 14: Chain of cells. The total magnetic flux density is computed as $\boldsymbol{b}=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{b}_{k}$. Arbitrary combinations of cells can be put in series. The represented arrangement is relevant for a multifilamentary strand, as explained in Section 7.
The advantage of this approach is that each cell can be solved independently, as they are all subject to the same magnetic field $\boldsymbol{h}$. Fixed point iterations can be performed to account for field-dependent parameters $\kappa$ or $\chi$ that should depend on the total magnetic flux density $\boldsymbol{b}$. The equations for each cell are identical to what was described in Sections 3.2 and 3.3, but each cell only contributes partly to the total magnetic flux density $\boldsymbol{b}$, with a fraction $\alpha_{k}\boldsymbol{b}_{k}$.
Once parameters are identified (see Section 5), the chain of cells defines the hysteresis model, i.e., the constitutive relationship $\boldsymbol{b}=\mathcal{B}(\boldsymbol{h})$ that reproduces the composite strand magnetization and loss.
4 Implementation and Inclusion in FE Framework
The hysteresis model output the magnetic flux density $\boldsymbol{b}(t)$ as a function of a given magnetic field $\boldsymbol{h}(t)$ (from which we can compute the magnetization $\boldsymbol{m}=\boldsymbol{b}/\mu_{0}-\boldsymbol{h}$ and the power loss). Both $\boldsymbol{b}(t)$ and $\boldsymbol{h}(t)$ are vector quantities varying in time. For a numerical simulation, time is discretized and the solutions are sought at successive time steps, based on the knowledge of the solution at previous time steps and on the new value of the magnetic field.
In this section, we first propose update rules for the two types of cells defined in the previous section, the S cell and CS cell. We then discuss the case of a chain of cells and conclude by describing how to include the hysteresis model in a FE framework.
4.1 S cell
If the irreversibility parameter $\kappa$ is constant, the S cell can be solved with an explicit update rule, as proposed in [27]. Let $\boldsymbol{h}_{\text{rev(p)}}$ be the reversible field computed at the previous time step. For a magnetic field $\boldsymbol{h}$, the updated reversible field $\boldsymbol{h}_{\text{rev}}$ reads:
$\boldsymbol{h}_{\text{rev}}$ =$&\boldsymbol{h}_{\text{rev(p)}},&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},&&\text{otherwise,}\end{aligned}\right.$
$=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}}).$ (24)
The associated magnetic flux density $\boldsymbol{b}$ is directly given by Eq. (9), and the irreversible field $\boldsymbol{h}_{\text{irr}}$ can be deduced from Eq. (8). The instantaneous rates of stored and dissipated energies, $p_{\text{rev}}$ and $p_{\text{irr}}$, are readily obtained by Eqs. (12) and (13), respectively.
If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ as proposed in Section 3.4 to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit. In such a case however, we observed that a good approximate solution is found easily in a few fixed point iterations, that is, by solving successively Eq. (4.1) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
It is remarkable that the implementation of the S cell entirely consists in the simple test of Eq. (4.1). It only requires the knowledge of the previous reversible field, which has to be stored as an internal variable in the numerical implementation.
Note that by contrast with ferromagnetic hysteresis with a nonlinear saturation law [27, 29], in this context of superconducting hysteresis, the relationship Eq. (9) is linear. The update rule Eq. (4.1) is therefore exact [29], up to the time discretization error.
4.2 CS cell
For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. We define
$\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}},$ (25)
and denote by $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
$\boldsymbol{g}=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (26)
with the update rule defined by Eq. (4.1) but as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$. This operation contains the first test, after which we can already compute the irreversible field $\boldsymbol{h}_{\text{irr}}=\boldsymbol{h}-\boldsymbol{g}$.
We proceed with the second test, encoded in Eq. (18). We first assume that $\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\leq\chi$ is valid (and we will correct the assumption if it is not the case). If $\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\leq\chi$, then we have the trial fields
$\boldsymbol{g}$ $=\boldsymbol{h}_{\text{rev}}^{\text{trial}}+\boldsymbol{h}_{\text{eddy}}^{\text{trial}}+\boldsymbol{h}_{\text{coupling}}^{\text{trial}}$ (27)
$=\boldsymbol{h}_{\text{rev}}^{\text{trial}}+\tau_{\text{e}}\dot{\boldsymbol{h}}_{\text{rev}}^{\text{trial}}+\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}^{\text{trial}}$ (28)
$\approx\boldsymbol{h}_{\text{rev}}^{\text{trial}}+\frac{\tau_{\text{e}}+\tau_{\text{c}}}{\Delta t}(\boldsymbol{h}_{\text{rev}}^{\text{trial}}-\boldsymbol{h}_{\text{rev(p)}}),$ (29)
with $\boldsymbol{h}_{\text{rev(p)}}$ the reversible field at the previous time step and $\Delta t$ the current time step. Solving this equation for $\boldsymbol{h}_{\text{rev}}^{\text{trial}}$, and then evaluating $\boldsymbol{h}_{\text{coupling}}^{\text{trial}}=\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}^{\text{trial}}$ gives the trial value (we drop the $\approx$ sign)
$\boldsymbol{h}_{\text{coupling}}^{\text{trial}}=\frac{\tau_{\text{e}}}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\left(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\right).$ (30)
If the assumption $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|\leq\chi$ is indeed satisfied, then:
$\boldsymbol{h}_{\text{coupling}}$ $=\boldsymbol{h}_{\text{coupling}}^{\text{trial}},$ (31)
$\boldsymbol{h}_{\text{eddy}}$ $=\frac{\tau_{\text{c}}}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\left(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\right).$ (32)
Otherwise, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, we have instead (Eq. (18)):
$\boldsymbol{h}_{\text{coupling}}$ $=\chi\frac{\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|},$ (33)
$\boldsymbol{h}_{\text{eddy}}$ $=\frac{\tau_{\text{c}}}{\Delta t+\tau_{\text{c}}}\left(\boldsymbol{g}-\boldsymbol{h}_{\text{coupling}}-\boldsymbol{h}_{\text{rev(p)}}\right).$ (34)
In both cases, the reversible field is obtained via
$\boldsymbol{h}_{\text{rev}}$ $=\boldsymbol{g}-\boldsymbol{h}_{\text{coupling}}-\boldsymbol{h}_{\text{eddy}}.$ (35)
All instantaneous power quantities can be computed by Eqs. (12), (13), (17), (20), and (21).
In the case of field-dependent irreversibility parameters $\kappa$ and $\chi$, a good convergence is again obtained in a few fixed points iterations, as for the S cell.
The simulation of the CS cell requires to save the values of vectors $\boldsymbol{g}$ and $\boldsymbol{h}_{\text{rev}}$ at each time step.
4.3 Chain of cells
In a hysteresis model with a chain of cells as proposed in Section 3.5, each cell $k$ is subject to the same magnetic field $\boldsymbol{h}$ and produces a distinct magnetic flux density $\boldsymbol{b}_{k}$. The total magnetic flux density $\boldsymbol{b}$ is computed as a weighted sum of the $\boldsymbol{b}_{k}$ as defined in Eq. (23). This is the only additional step compared to the single cell case.
Apart from this change, each cell can be solved independently exactly as described above, possibly with global fixed point iterations in the case of field-dependent irreversibility parameters $\kappa$ and $\chi$ that depend on the total magnetic flux density $\boldsymbol{b}$.
4.4 Inclusion in a finite element $\phi$-formulation
[…]
Let us consider a numerical domain $\Omega$. It is decomposed into the superconducting strand (or any other superconducting system) to homogenize, denoted as $\Omega_{\text{m}}$, and the complementary domain, denoted as $\Omega_{\text{m}}^{\text{C}}$. For simplicity, we assume that $\Omega_{\text{m}}^{\text{C}}$ is non-conducting, but the approach can be easily extended to conducting domains as well. Also, in this paper, we only consider cases with no transport current in $\Omega_{\text{m}}$.
The set of equations to solve is therefore:
[…]
These equations can be solved numerically with the finite element method, on a mesh, i.e., a spatial discretization of the numerical domain $\Omega$.
As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is more natural to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [41, 42]. In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, this formulation can be reduced to a $\phi$-formulation [29].
Starting from an initial solution, it consists of finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
[…]
with $\Phi(\Omega)$ and $\Phi_{0}(\Omega)$ appropriate function spaces, and where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$.
The presence of the hysteresis law in Eq. (37) makes the system history-dependent. The fields $\boldsymbol{h}_{\text{rev}}$ and $\boldsymbol{g}$ must therefore be saved at each steps. They are defined in $\Omega_{\text{m}}$ only and are chosen to be element-wise constant vectors.
The hysteresis law also makes the system nonlinear. An iterative scheme is therefore necessary for a numerical simulation. In this case, we observed that the Newton-Raphson method leads to efficient simulations. From an initial iterate $\boldsymbol{h}_{0}$, it consists in solving successively a linearized version of Eq. (37) until a given convergence criterion is met. At iteration $i\in\mathbb{N}_{0}$, the formulation reads, in terms of the unknown field $\boldsymbol{h}_{i}$,
[…]
$+\left(\mu_{0}\boldsymbol{h}_{i}\,,\boldsymbol{h}^{\prime}\right)_{\Omega_{\text{m}}^{\text{C}}}=0,$ (38)
[…]
5 Parameter Identification
The hysteresis model presented in Section 3 provides a flexible tool able to capture a variety of different hysteretic responses thanks to the approach with several cells connected as a chain. The parameter values of each of these cells must be properly chosen in order to reproduce reference solutions or measurement data.
In this section, we propose a simple identification procedure for chains of S cells and explain a heuristic for chains of CS cells, which will be further illustrated in Section 7. The reference solutions are obtained by detailed numerical models such as those presented in Section 2.
5.1 Chain of S cells
In a chain of S cells, there are $2N$ parameters to be defined: the constant weights $\alpha_{k}$ and the irreversibility parameters $\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa}_{k}$. The number of cells $N$ must also be chosen as a trade-off between computational cost, implementation easiness and accuracy.
As a reference solution, we consider the magnetization curve of a superconducting strand subject to a unidirectional field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{app}}(t)\hat{\boldsymbol{e}}_{y}=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$, with a sufficiently large field $h_{\text{max}}$ for the system to be fully penetrated, and at a sufficiently low frequency $f$ for the coupling and eddy current effects to be negligible. We compute the associated average magnetic flux density in the strand $\boldsymbol{b}_{\text{in}}$ as
$\boldsymbol{b}_{\text{in}}$ $=\mu_{0}\boldsymbol{h}_{\text{in}}+\mu_{0}\boldsymbol{m}=\mu_{0}\left(\boldsymbol{h}_{\text{app}}+\frac{1}{2}\boldsymbol{m}\right),$ (39)
using Eqs. (1) and (7). An example of such a curve for the composite strand described in Section 2 is given in Fig. 15, where $\mu_{0}h_{\text{max}}=2$ T. The curve is obtained with a standard 2D $h$-$\phi$-formulation that does not model coupling currents between the filaments, at $f=0.01$ Hz.
Figure 15: Reference magnetization curve for the identification of a chain of S cells.
[…]
Step 1: Fitting the field-dependent scaling
The first step is to fit the irreversibility parameter scaling, $f_{\kappa}(\boldsymbol{b})$, introduced in Eq. (22) to account for the field-dependent critical current density. For this, we consider parts of the magnetization cycle in which the superconductor is fully magnetized. In Fig. 15, this corresponds to the purple curves labelled as Step 1.
For a chain of S cell, using Eqs. (8) and (23), the total magnetization is given by
$\boldsymbol{m}$ $=\boldsymbol{b}/\mu_{0}-\boldsymbol{h}=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{h}_{\text{rev},k}-\boldsymbol{h}$
$=\sum_{k=1}^{N}\alpha_{k}\left(\boldsymbol{h}_{\text{rev},k}-\boldsymbol{h}\right)=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{h}_{\text{irr},k}.$ (40)
In the fully magnetized situation and with a unidirectional excitation, we have
$|m|=\sum_{k=1}^{N}\alpha_{k}\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\underbrace{\sum_{k=1}^{N}\alpha_{k}\bar{\kappa}_{k}}_{m_{\text{max}}},$ (41)
with $m_{\text{max}}$ (in A/m) the maximum magnetization. As a result, the scaling function $f_{\kappa}(\boldsymbol{b})$ directly describes the shape of the purple curves in Fig. 15.
A simple approach consists in choosing $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ as a scaling over the whole field range. This however does not lead to very accurate results at low fields for which $j_{\text{c}}$ strongly depends on $\boldsymbol{b}$. The reason is that the strand magnetization depends on the actual field distribution inside it, which is not uniform, and not only on the average vector $\boldsymbol{b}_{\text{in}}$. The larger the filaments in the strand, the larger the variation of $\boldsymbol{b}$ within each filament.
A better approach is to directly identify the scaling $f_{\kappa}(\boldsymbol{b})$ on the reference curve. The purple curves can therefore be used to directly define $f_{\kappa}(\boldsymbol{b})$ in the available field range. Outside of the field range, $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ can be used as a first approximation.
In some circumstances, the maximum magnetization may not be observed at $\boldsymbol{b}_{\text{in}}=\boldsymbol{0}$ exactly. In such a case, the scaling $f_{\kappa}$ can be written in terms of a weighted sum such as $u\boldsymbol{b}+(1-u)\mu_{0}\boldsymbol{h}$, with $u\in[0,1]$ to be chosen.
Step 2: Choosing the $\bar{\kappa}_{k}$ and fitting the weights $\alpha_{k}$
The second step consists in finding the remaining parameters, $\bar{\kappa}_{k}$ and $\alpha_{k}$, in order to reproduce transition branches between fully magnetized states, such as those labelled as Step 2 in Fig. 15.
The proposed procedure consists in choosing a priori the $N$ values $\bar{\kappa}_{k}$ and then fixing the weights $\alpha_{k}$ accordingly. For simplicity, we arrange the $\bar{\kappa}_{k}$ values in increasing order with respect to $k$ and start with $\bar{\kappa}_{1}=0$ A/m.
We consider the virgin magnetization curve represented by the dotted curve in Fig. 15. Initially, $h_{\text{rev},k}=0$, $\forall k=1,\dots,N$. As the magnetic field $h_{\text{in}}$ progressively increases, it reaches the surface of the cell spheres one by one (see Fig. 8). This happens at successive threshold field values that depend on the $\bar{\kappa}_{k}$ constants and on the $f_{\kappa}(b_{\text{in}})$ scaling determined in Step 1. We denote these threshold fields as $h_{k}$.
For $h<h_{k}$, we have $h_{\text{rev},j}=0$, $\forall j=k,\dots,N$, such that only the first $k-1$ cells contribute to the magnetic flux density $b$. The weights $\alpha_{k-1}$ can therefore be successively determined by forcing the solution of the hysteresis model to match the reference solution at each $h_{k}$, for $k=2,\dots,N$. The last weight $\alpha_{N}$ is calculated so that all weights add up to one.
The choice of the number of cells $N$ and the a priori distribution of values for the $\bar{\kappa}_{k}$ depend on the sought accuracy and application. This is discussed in Section 6.
5.2 Chain of CS cells
For a chain of CS cells, there are $5N$ parameters to choose: $\alpha_{k}$, $\kappa_{k}$, $\tau_{\text{e},k}$, $\tau_{\text{c},k}$, and $\chi_{k}$, for $k=1,\dots,N$. To identify all these parameters, reference solutions at different frequencies are necessary.
A reference solution at a sufficiently low frequency for the eddy and coupling current effects to be negligible can be used exactly as described in the previous section to identify the $\alpha_{k}$ and $\kappa_{k}$ (with both the $f_{\kappa}(\boldsymbol{b})$ scaling and the $\bar{\kappa}_{k}$ values). This leads to a number of $3N$ remaining parameters.
The time constants $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ can be chosen so that the positions of the peaks in eddy and coupling losses respectively correspond to those of reference solutions, such as shown in Fig. 3. Choosing identical values for all the cells already leads to good results within limited amplitude ranges. To reproduce the observation that the maxima in coupling and eddy loss are observed at lower frequencies when the field amplitude increases, one can choose larger values of $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ for cells associated with larger irreversibility parameters $\kappa_{k}$. This will be illustrated in Section 7.
Once the time constants are chosen, the only remaining parameters are the irreversibility parameters $\chi_{k}(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi}_{k}$. The scaling $f_{\chi}(\boldsymbol{b})$ can be chosen as in the static case (Step 1), based on a reference solution at a sufficiently high frequency for the dynamic effects to be visible in the magnetization curve, e.g., at $f=100$ Hz for the 54-filament strand, as shown in Fig. 7. Then, the values $\bar{\chi}_{k}$ can be identified in order to best reproduce the transition branches, as in Step 2 of the static case, but now with values $\bar{\chi}_{k}$ as unknowns instead of the weights $\alpha_{k}$ which are already fixed.
6 Results with a chain of S cells
(a) Major loop $h_{\text{app}}^{(1)}(t)$.
(b) Minor loops $h_{\text{app}}^{(2)}(t)$.
Figure 16: Comparison of the hysteresis model results with a reference solution. Unidirectional transverse magnetic field. Chain of S cells with $N=6$ cells, parameters of Table 2. (a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$. (b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$. (Up) Magnetization loop. (Middle) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section. (Down) Internal magnetic field, the input of the hysteresis model.
The first application consists in the 54-filament strand presented in Section 2, subject to a transverse field varying sufficiently slowly for the eddy and coupling current effects to be negligible, that is, with frequencies of the order of $f=0.01$ Hz. For the reference solution, we assume a non-conducting matrix and use a classical $h$-$\phi$-formulation [41], so that coupling effects are completely removed.
Because of the finite value of the power index $n=30$, the strand response is not truly rate-independent (this would only be the case at the limit $n\to\infty$). Still, we model the macroscopic strand response (magnetization and loss) with a chain of S cells as a first approximation, and show that it already provides very good results.
We start the analysis by identifying the parameters of the chain of S cells. We then compare the prediction of the resulting model for different types of excitations, both unidirectional and bidirectional.
6.1 Parameter identification
The reference solution is the major loop represented in Fig. 15. We choose $N=6$ and apply the two-step procedure described in Section 5.1 with $N$ equally spaced values $\mu_{0}\bar{\kappa}_{k}$ from 00 T to $0.75$ T. The obtained parameters are given in Table 2.
$k$ $\alpha_{k}$ $\mu_{0}\bar{\kappa}_{k}$
[…]
$1$ $48.24$ 00
$2$ $18.07$ $150$
$3$ $16.99$ $300$
$4$ $9.31$ $450$
$5$ $5.16$ $600$
$6$ $2.22$ $750$
Table 2: Parameters of a chain of S cells with $N=6$, identified on the reference magnetization loop shown in Fig. 15.
It is interesting to notice that the first S cell, which is anhysteretic, almost contributes to half of the magnetic flux density. This is due to the large fraction of normal conductor in the strand, the matrix, which does not behave as a hysteretic material.
6.2 Unidirectional excitation
Using the parameters obtained above, we compare the predictions of the hysteresis model in terms of magnetization and loss for two unidirectional excitations along $\hat{\boldsymbol{e}}_{y}$. The first one is a harmonic field:
$\mu_{0}h_{\text{app}}^{(1)}(t)$ $=\mu_{0}h_{\text{max}}\sin(2\pi ft),$ (42)
with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. The second one is a biharmonic field, so as to produce minor magnetization loops:
$\mu_{0}h_{\text{app}}^{(2)}(t)$ $=\mu_{0}h_{\text{max}}\left(\sin(\pi ft)+0.25\sin(6\pi ft)\right).$ (43)
Results are given in Fig. 16. With $N=6$ cells, the hysteresis model is capable of reproducing faithfully both major and minor magnetization loops. The loss is accurately reproduced in both cases, with a relative error on the total loss below $1\%$.
The influence of the number of cells on the relative error on total loss is shown in Fig. 17. For the considered excitations, increasing the number of cells helps reducing the error for low $N$ values ($\leq 3$), but the error then stabilizes to a non-zero value. One source of error is related to the scaling function $f_{\kappa}(\boldsymbol{b})$ for irreversibility parameters, which is only approximate. Indeed, in reality, the magnetic flux density $\boldsymbol{b}$ is not uniform in the filament, but rather varies due to the screening currents and geometrical effects. Also, as mentioned above, the reference solution is not truly rate-independent, due to the finite $n$ value in the power law.
Figure 17: Evolution of the relative error $(Q_{\text{hyst}}-Q_{\text{ref}})/Q_{\text{ref}}$ between total losses predicted by the hysteresis model ($Q_{\text{hyst}}$) and the reference model ($Q_{\text{ref}}$), for different sources, as a function of the number of cells $N$, with equally spaced values $\mu_{0}\bar{\kappa}_{k}$, from 00 T to $0.75$ T. In the case $N=1$, $\mu_{0}\bar{\kappa}_{1}=0.168$ T and $\alpha_{1}=1$.
The major advantage of the method is its computational efficiency. Once the model parameters are identified, solving the hysteresis model is tremendously faster than solving a finite element model (typically a few milliseconds compared to a few tens of minutes). This efficiency is crucial in view of homogenizing full-scale superconducting systems.
In practice, the number of cells and the distribution of $\bar{\kappa}_{k}$ values must be chosen depending on the actual excitations to be considered in the end application. A chain of S cells produces no loss for field variations smaller than the smallest non-zero irreversibility parameter $\kappa_{k}(\boldsymbol{b})$, as illustrated in Fig. 18. To reproduce accurately power loss at low fields, or for low field ripples, the chain of S cells must therefore contain cells with sufficiently small irreversibility parameters. This will be further illustrated in the case of a chain of CS cells in Section 7.
Figure 18: Illustration of a chain of S cells with $N=3$. The centers of the spheres are at positions $\boldsymbol{h}_{\text{rev},k}$, established by the history of $\boldsymbol{h}$. Variations of the magnetic field $\boldsymbol{h}$ that stay inside the smallest sphere do not produce any loss.
6.3 Bidirectional excitation
The parameters were identified based on the results of a unidirectional excitation, but all the hysteresis model equations are vectorial, and the model is therefore directly applicable to general excitations. We show in Fig. 19 the results for a bidirectional rotating transverse field excitation defined by:
$\mu_{0}\boldsymbol{h}_{\text{app}}^{(3)}(t)=\quad$ $\mu_{0}h_{\text{max}}\sin(\pi ft)\sin(\pi ft)\hat{\boldsymbol{e}}_{x}$
$+$ $\mu_{0}h_{\text{max}}\sin(\pi ft)\cos(\pi ft)\hat{\boldsymbol{e}}_{y},$ (44)
with $\mu_{0}h_{\text{max}}=2$ T, $f=0.01$ Hz, and $\hat{\boldsymbol{e}}_{x}$, $\hat{\boldsymbol{e}}_{y}$ the unit vectors in $x$-, $y$-directions. The influence of the number of cells $N$ on the relative error is shown in Fig. 17, a constant error of $\approx 3\%$ remains for large values of $N$.
Figure 19: Comparison of hysteresis model results with a reference solution. Rotating transverse magnetic field $\boldsymbol{h}_{\text{app}}^{(3)}(t)$. Chain of S cells with $N=6$ cells and parameters of Table 2. Solutions at selected time instants are indicated by the colored circles (reference solution) and crosses (hysteresis model). (Up) Internal magnetic field and magnetization of the filament (same legend as in the bottom sub-figure). Values are in tesla (T). (Down) Dissipated power per unit length as a function of time.
The hysteresis model reproduces well the magnetization and power loss. The angle between the inner field $\boldsymbol{h}_{\text{in}}$ and the magnetization vector $\boldsymbol{m}$ is faithfully described, as shown by the solutions at selected time instants in Fig. 19. The agreement is not perfect, but very satisfying provided that the model parameters were fully identified using results from a unidirectional situation, with absolutely no information about the strand response in more general cases. We let the questions of improving the agreement and tackling anisotropic systems for further works.
7 Results with a chain of CS cells
As a generalization of the previous section, we consider the same 54-filament strand subject to a transverse field, but now with higher rates of field changes, so that both magnetization and loss exhibit frequency dependence. The reference solutions are obtained as described in Section 2 with the CATI method in order to account for coupling current effects.
We start the analysis by choosing the hysteresis model parameters based on reference solutions. We then illustrate the results for wide ranges of field amplitudes and frequencies.
7.1 Parameter identification
We choose the chain structure represented in Fig. 14. Compared to the chain of S cells of the previous section, two cells now have a zero irreversibility parameter, $\kappa_{1}=0$ A/m and $\kappa_{2}=0$ A/m. They allow to reproduce losses at low fields, for which filament loss is small with respect to coupling and eddy loss. These two cells allow to reproduce the transitions between the three regimes illustrated in Fig. 20 at low fields, associated with different magnetization slopes, as shown in Fig. 21. The curves are obtained with the parameter values of Table 3, which are discussed below.
Figure 20: Field distribution at low fields ($\mu_{0}h=10$ mT here) and different frequencies. The lines are parallel to the local magnetic field, the larger and darker they are, the higher the field amplitude. (a) Uncoupled filament and limited magnetization. (b) Increased magnetization due to coupled filaments. (c) Close to perfect diamagnetism due to eddy currents.
Figure 21: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity). Dashed lines are reference solutions and solid lines are from the hysteresis model.
At low frequencies, the dominant contribution to magnetization and loss comes from superconducting filament hysteresis. We can reuse the material parameters found in the previous section for the chain of S cells as was given in Table 2. In order to reproduce losses at fields lower than $0.15$ T, we also introduce $8$ new cells with lower irreversibility parameters, logarithmically spaced from $1$ mT to $80$ mT. The associated weights can be identified as proposed in Section 5.1, or based on analytical solutions at low fields.
We then choose the time constants. A good fit with the reference solution is obtained with the values provided in Table 3. Larger time constants are chosen for cells associated with higher irreversibility parameters, in order to reproduce the shifts of the peak coupling and eddy loss observed in Fig. 3 and discussed in Section 2.3.
Finally, we fix the scaling $f_{\chi}(\boldsymbol{b})$ and the parameters $\bar{\chi}_{k}$. We found that a decent fit is obtained with
$f_{\chi}(\boldsymbol{b})=\frac{1-\|\boldsymbol{b}\|/b_{0}}{1+\|\boldsymbol{b}\|/b_{1}},$ (45)
with $b_{0}=15$ T and $b_{1}=4$ T. The parameters $\bar{\chi}_{k}$ are then tuned manually to obtain the values in Table 3.
$k$ $\alpha_{k}$ $\mu_{0}\bar{\kappa}_{k}$ $\tau_{\text{e},k}$ $\tau_{\text{c},k}$ $\mu_{0}\bar{\chi}_{k}$
[…]
$1$ $16.38$ $0.0$ $0.02$ $0.00$ $0.00$
$2$ $22.62$ $0.0$ $0.02$ $0.08$ $1.13$
$3$ $0.05$ $1.0$ $0.02$ $0.08$ $0.01$
$4$ $0.10$ $1.9$ $0.02$ $0.08$ $0.02$
$5$ $0.19$ $3.5$ $0.02$ $0.08$ $0.03$
$6$ $0.35$ $6.5$ $0.02$ $0.08$ $0.07$
$7$ $0.66$ $12.2$ $0.02$ $0.08$ $0.09$
$8$ $1.24$ $22.9$ $0.02$ $0.08$ $0.14$
$9$ $2.32$ $42.9$ $0.02$ $0.08$ $0.21$
$10$ $4.33$ $80.2$ $0.02$ $0.08$ $0.40$
$11$ $18.07$ $150.0$ $0.30$ $0.28$ $1.50$
$12$ $16.99$ $300.0$ $0.30$ $0.28$ $2.10$
$13$ $9.31$ $450.0$ $0.30$ $0.28$ $3.15$
$14$ $5.16$ $600.0$ $0.30$ $0.28$ $4.20$
$15$ $2.22$ $750.0$ $0.30$ $0.28$ $5.25$
Table 3: Parameters of a chain of CS cells with $N=15$.
7.2 Results and numerical performance
Using the model parameters given in Table 3, we compare the predictions of the hysteresis model with the reference solutions. The power loss per cycle as a function of frequency and selected magnetization loops are given in Figs. 22 and 23. The power loss per cycle is also presented as a map in Fig. 24, by interpolating the results of $400$ simulations, spanning over $25$ frequencies and $16$ field amplitudes, logarithmically spaced.
Figure 22: Power loss per cycle for three different field amplitudes as a function of frequency. Solid curves are results from the hysteresis model with parameters of Table 3 and dashed curves are reference solutions. The legend is the same for the three subfigures.
Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies. Solid curves are results from the hysteresis model with parameters of Table 3 and dashed curves are reference solutions.
Figure 24: Total loss per cycle and per unit length as a function of the frequency and amplitude of the applied field. Dashed curves are contour lines of equal loss. Solid curves delimit areas of distinct dominant loss contributions (from left to right: hysteresis in uncoupled filaments, coupling loss, hysteresis in coupled filaments, and eddy losses). (Left) Reference model, adapted from [31]. (Right) Prediction of the hysteresis model with the parameters given in Table 3.
The overall agreement on the total loss and magnetization is very good. The hysteresis model correctly reproduces the different regimes of distinct dominant loss contributions and describes the magnetization faithfully. The difference with the reference solution is higher at high frequency and high fields, as can be seen on the contour lines in the top right corners in Fig. 24. This can be improved with a careful choice of model parameters, but as discussed in Section 3.3, the matching will never be perfect in the limit $f\to\infty$ as the CS cells do not reproduce a $1/\sqrt{f}$ decrease of the loss per cycle, but rather a $1/f$ decrease. For not too high frequencies however, the importance of this effect is limited.
The major advantage of the hysteresis model compared to the detailed simulations is the very small associated computational work. Even with $N=15$ cells, the whole loss map on the right of Fig. 24, which consists of $400$ hysteresis model simulations, can be performed in a few seconds, whereas the reference model requires $26$ days of computation time in total to generate the loss map on the left of Fig. 24.
[…]
Figure 25: Solutions of the reference CATI model and the ROHM model for $\mu_{0}h_{\text{max}}=1$ T and $f=1$ Hz. (a-b) Solutions at $t=0.25$ s. (c-d) Solutions at $t=0.5$ s. Arrows outside of the strand represent the magnetic flux density $\boldsymbol{b}$ (shared color scale in the middle). Colored elements in (a) and (c) represent the $z$-component of the current density $\boldsymbol{j}$ (color scale on the left, with $j_{\text{c0}}=3\times 10^{10}$ A/m2). Arrows inside the strand in (b) and (d) represent the magnetization $\mu_{0}\boldsymbol{m}=\boldsymbol{b}-\mu_{0}\boldsymbol{h}$ (color scale on the right).
In superconducting magnets, many strands are combined together into cables, which themselves are wound around the magnet aperture. The hysteresis model can be used to homogenize the large number of strands into an equivalent homogeneous material. In such a case, which is let for further work, the reference detailed solution should be that of a close-packed set of strands, which can typically be modelled via appropriate periodic boundary conditions [14], rather than a single strand in air, as was done here for the sake of illustration. This may result in different model parameters, but the general approach is unchanged.
8 Conclusion
In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous loss in composite superconductors subject to transient external fields. This model is designed by adapting to superconducting systems the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27]. We focused specifically on the case of a multifilamentary superconducting strand and proposed two models: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects.
We proposed a parameter identification approach for both models and described their inclusion in a finite element framework. Finally, we demonstrated that these models help to strongly reduce the computational cost compared to conventional simulations, while offering a very good accuracy in a wide range of field amplitudes and rates of field change.
This work constitutes a first step towards the homogenization of superconducting magnets, in which the large number of turns makes detailed simulations unrealistically expensive. Replacing the coil windings by homogenized material properties such as the proposed ROHM model allows to describe magnetization and loss efficiently. Further work is necessary to include non-zero transport current, which was not covered in this paper.
Extension of the method to anisotropic systems such as high-temperature superconducting tapes or stacks of tapes in view of their homogenization can also be considered as further work.
Appendix A Magnetization in 2D Problems
In this section, we justify the introduction of a factor $2$ in the magnetization Eq. (1), in the case of an infinitely long problem solved in 2D with perpendicular currents.
The magnetic dipole moment $\boldsymbol{M}$ of a current density distribution $\boldsymbol{j}$ in a given volume $V$ is evaluated as
$\boldsymbol{M}=\frac{1}{2}\int_{V}\boldsymbol{x}\times\boldsymbol{j}\,\text{d}V,$ (46)
with $\boldsymbol{x}$ the position vector. Consider a rectangular current tube of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig. 26. Accounting for the four sides of the rectangle, the contribution of this current loop to the total magnetic dipole moment is $wl\,\text{d}I$. Long and short sides of the rectangle contribute equally to the total magnetic dipole moment.
Figure 26: Magnetic dipole moment of a current tube in a cylinder.
When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D and per unit length, on a cross section of the conducting cylinder. In such a case, however, one must still account for the fact that all current loops close at the end of the cylinder, i.e., at infinity, for computing the magnetization per unit length. These closing currents are not part of the 2D solution, and are then missing from any magnetization calculation performed solely on the 2D cross section. Because the closing currents contribute to exactly half of the magnetic dipole moment, one can account for them by introducing a factor $2$, as is done in Eq. (1).
Appendix B Applied Field and Internal Field
The power loss per cycle $Q_{\text{tot}}$ given by Eq. (4) can equivalently be computed in terms of $\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\boldsymbol{m}/2$. Indeed, as also shown in [38],
$Q_{\text{tot}}$ $=a\oint\mu_{0}\boldsymbol{h}_{\text{app}}\cdot\text{d}\boldsymbol{m}$
$=a\oint\mu_{0}(\boldsymbol{h}_{\text{in}}+\boldsymbol{m}/2)\cdot\text{d}\boldsymbol{m}$
$=a\oint\mu_{0}\boldsymbol{h}_{\text{in}}\cdot\text{d}\boldsymbol{m}+\frac{a\mu_{0}}{2}\underbrace{\oint\boldsymbol{m}\cdot\text{d}\boldsymbol{m}}_{=0\text{ if closed}}$
$=a\oint\mu_{0}\boldsymbol{h}_{\text{in}}\cdot\text{d}\boldsymbol{m}.$ (47)
That is, the areas enclosed by the loops represented in Figs. 5 and 7 are identical.
Appendix C Hysteresis Model Jacobian
[…]
$\frac{\partial\boldsymbol{b}}{\partial\boldsymbol{h}}=\sum_{k=1}^{N}\alpha_{k}\mu_{0}\frac{\partial\boldsymbol{h}_{\text{rev},k}}{\partial\boldsymbol{h}}.$ (48)
Its expression can be computed independently for each cell. Below, we give its expression for the simple case of a S cell and then generalize for a CS cell. We drop the index $k$ for conciseness.
C.1 S cell
For a S cell, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated with Eq. (4.1). The derivative of this update rule is given by
$\frac{\partial\boldsymbol{h}_{\text{rev}}}{\partial\boldsymbol{h}}$ =$&\mathbf{0},\quad\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\left(1-\frac{\kappa}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|}\right)\mathbf{I}+\frac{\kappa}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|^{3}}\mathbf{K},\\ &\qquad\text{otherwise,}\end{aligned}\right.$
$=\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}}),$ (49)
with $\mathbf{I}$ the second-order identity tensor and $\mathbf{K}$ defined by the following dyadic product:
$\mathbf{K}=(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}).$ (50)
C.2 CS cell
For a CS cell, we first compute $\boldsymbol{h}_{\text{coupling}}^{\text{trial}}$ using Eq. (30). In the case $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|\leq\chi$, using Eqs. (35), (31), and (32), we have the reversible field given by
$\boldsymbol{h}_{\text{rev}}=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\left(\boldsymbol{g}+\frac{\tau_{\text{e}}+\tau_{\text{c}}}{\Delta t}\boldsymbol{h}_{\text{rev(p)}}\right)$ (51)
whose derivative reads
$\frac{\partial\boldsymbol{h}_{\text{rev}}}{\partial\boldsymbol{h}}$ $=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\frac{\partial\boldsymbol{g}}{\partial\boldsymbol{h}}$
$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (52)
where $\boldsymbol{\mathcal{W}}_{\kappa}$ is the second-order tensor defined in Eq. (C.1), written here as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$.
By contrast, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, one must use Eqs. (35), (33), and (34), which yields the following reversible field
$\boldsymbol{h}_{\text{rev}}=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\left(\boldsymbol{g}-\chi\frac{\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|}+\frac{\tau_{\text{e}}}{\Delta t}\boldsymbol{h}_{\text{rev(p)}}\right)$ (53)
whose derivative reads
$\frac{\partial\boldsymbol{h}_{\text{rev}}}{\partial\boldsymbol{h}}$ $=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\left(\frac{\partial\boldsymbol{g}}{\partial\boldsymbol{h}}-\frac{\partial}{\partial\boldsymbol{g}}\left(\chi\frac{\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|}\right)\cdot\frac{\partial\boldsymbol{g}}{\partial\boldsymbol{h}}\right)$
$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\boldsymbol{\mathcal{W}}_{\chi}(\boldsymbol{g},\boldsymbol{h}_{\text{rev(p)}})\cdot\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (54)
where $\boldsymbol{\mathcal{W}}_{\chi}$ is the second-order tensor defined by
$\boldsymbol{\mathcal{W}}_{\chi}(\boldsymbol{g},\boldsymbol{h}_{\text{rev(p)}})=\left(1-\frac{\chi}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|}\right)\mathbf{I}$
$+\frac{\chi}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|^{3}}\mathbf{L},$ (55)
with
$\mathbf{L}=(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}).$ (56)

### v2

@@
 Abstract
-In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors.
-Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without having to solve for the detailed current density distribution.
-It can be used as a homogenized model in large-scale superconducting systems in order to significantly reduce the computational work compared to detailed simulations.
-In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to an external transverse magnetic field.
-We propose two levels of ROHM models: a rate-independent model that reproduces pure hysteresis without dynamic effects, and a rate-dependent model that generalizes the former by also reproducing dynamic effects observed in superconducting strands due to coupling and eddy currents.
-We describe the implementation and inclusion of these models in a finite element framework, discuss their parameter identification and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency.
-  • September 20, 2024
+In this paper, we propose the Reduced Order Hysteretic Magnetization (ROHM) model to describe the magnetization and instantaneous power loss of composite superconductors subject to time-varying magnetic fields.
+Once the parameters of the ROHM model are fixed based on reference simulations, it allows to directly compute the macroscopic response of composite superconductors without the need to solve the detailed current density distribution.
+It can then be used as part of a homogenization method in large-scale superconducting models to significantly reduce the computational effort compared to detailed simulations.
+In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to a time-varying transverse magnetic field.
+We propose two levels of ROHM models: (i) a rate-independent model that reproduces hysteresis in the filaments, and (ii) a rate-dependent model that generalizes the first level by also reproducing dynamic effects due to coupling and eddy currents.
+We describe the implementation and inclusion of these models in a finite element framework, discuss how to deduce the model parameters, and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency over a wide range of excitation frequencies and amplitudes.
+  • October 25, 2024
 Keywords: Reduced Order Method, Hysteresis Model, AC Loss, Magnetization.
 […]
-*                                    This work is a preliminary version. Copyright may be transferred without notice, after which this version may no longer be accessible.
+*                                    This work has been submitted to a journal for possible publication.
+Copyright may be transferred without notice, after which this version may no longer be accessible.
 […]
 […]
-*                                    This work is a preliminary version. Copyright may be transferred without notice, after which this version may no longer be accessible.
+*                                    This work has been submitted to a journal for possible publication.
+Copyright may be transferred without notice, after which this version may no longer be accessible.
 1 Introduction
-Large-scale superconducting systems are multi-scale structures made of a high number of wounded superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes.
+Large-scale superconducting magnets are multi-scale structures made of a high number of wound superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes.
 The magnetic response of these systems is hysteretic and rate-dependent.
-Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting parts [1, 2, 3, 4], whereas rate-dependency is caused by eddy currents, and coupling current dynamics between different conductors [5, 6, 7].
+Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting elements [1, 2, 3, 4], whereas rate-dependency is caused by eddy and coupling current dynamics [5, 6, 7].
 Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8].
-Their computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
+Their accurate description and computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
 Most approaches to describe the superconducting hysteresis are based on the definition of a relationship between electric field and current density, either non-smooth, as in the Bean model [1], or smooth, as in the power law model [9].
@@
 Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes the numerical models computationally demanding to solve, already for single strands or cables [11].
-As a consequence, modelling large-scale superconducting systems in all their details down to the conductor level is completely unrealistic in practice.
-Faster, more efficient, methods are necessary.
+As a consequence, modelling large-scale superconducting systems in all their details down to the superconductor level is completely unrealistic in practice.
+A faster and more efficient method is necessary.
 Homogenization methods which reproduce the macroscopic effects of coupling dynamics without modelling them explicitly are good candidates.
 The general idea of these methods is to describe the systems in terms of the average fields [12], in order to reproduce the magnetization and loss accurately, without having to solve for the detailed current density distribution, which has the potential to strongly reduce the computational cost of the simulations.
-Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures [13, 14, 15] with bundles of wires of any shape, described in the frequency or time domain [16, 17].
-They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions, or on experimental measurement results.
-Following a similar approach for superconducting systems requires to use dedicated history-dependent parameters in order to reproduce their hysteretic response.
-A few hysteresis models have been proposed for modelling superconductor.
+Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures with bundles of wires of any shape [13, 14, 15], described in the frequency or time domain [16, 17].
+They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions.
+Following a similar approach for superconducting systems requires the use of dedicated history-dependent parameters in order to reproduce their hysteretic response.
+A few hysteresis models have been proposed for modelling superconductors.
 A vector model is proposed in [8, 18] for field quality computation in accelerator magnets.
@@
 First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions.
-Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that can be combined.
-Third, it is consistently defined as a vector, and not scalar, model from the beginning [27].
+Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that are combined to form the final model.
+Third, it is consistently defined as a vector, and not scalar model from the beginning [27], and hence applies to two-dimensional (2D) or three-dimensional (3D) settings.
 Fourth, it can be directly generalized to include rate-dependent effects [28].
-Finally, the equations of this model consist in simple explicit tests, which lead to straightforward implementations and very efficient simulations.
+Finally, the equations of this model consist in simple explicit tests, which leads to straightforward implementations and very efficient simulations.
 In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis by focussing on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis.
 This situation is relevant for most applications and constitutes a first step towards the homogenization of a complete magnet winding.
-This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects: eddy currents in normal conductors and interfilamentary coupling currents.
+This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects, eddy currents and interfilamentary coupling currents in the strand matrix, as well as filament hysteresis change due to coupling and eddy currents.
 For this paper, we assume that the strand carries no transport current.
@@
 We also consider a constant temperature and do not model thermal effects.
-We propose two models: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the first model by also describing the eddy current and coupling current effects, in addition to the superconducting hysteresis.
-The models are general and can be adapted to a variety of composite superconductors, such as strands, tapes, or cables.
-They can also be easily implemented in any FE framework allowing for user-defined material properties.
+We propose the Reduced Order Hysteretic Magnetization (ROHM) model, and present two variations of it: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the former by also describing eddy current and coupling current effects, in addition to the superconducting hysteresis.
+The ROHM model is general and can be adapted to a variety of superconductors or composite superconductors, such as strands, tapes, stacks of tapes, cables, or bulks.
+It can also be easily implemented in any FE framework that allows for user-defined material properties.
 This paper is structured as follows.
-In Section 2, we introduce the model of a multifilamentary strand and describe its response in terms of magnetization and loss for transverse fields in a range of frequencies and amplitudes.
-In Section 3, we present the equations leading to rate-independent and rate-dependent hysteresis models.
-In Section 4, we discuss the implementation of the equations and describe how to include them in a two-dimensional (2D) FE framework with a $\phi$-formulation.
-We then propose in Section 5 a parameter identification procedure for the models.
-Finally, in Sections 6 and 7, we apply the hysteresis models in rate-independent and rate-dependent situations, respectively, and demonstrate their potential in the context of homogenization techniques.
+In Section 2, we describe the response of a multifilamentary strand in terms of magnetization and loss for transverse fields with a range of frequencies and amplitudes, in order to motivate the design of the ROHM model.
+In Section 3, we present the equations defining the rate-independent and rate-dependent ROHM models.
+In Section 4, we discuss the implementation of these equations and we describe how to include them in a 2D FE framework with a $\phi$-formulation.
+We then propose in Section 5 a parameter identification approach for fixing the ROHM model parameters based on reference solutions.
+Finally, in Sections 6 and 7, we apply the ROHM model in rate-independent and rate-dependent situations, respectively, and demonstrate its potential in the context of homogenization techniques.
 2 Multifilamentary Strand Dynamics
-In this section, we present the response of a composite multifilamentary superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization.
-To this end, we model the detailed current density distribution inside the strand with a FE method and we analyze the different dynamics that result from the strand composite structure.
+In this section, we describe the response of a composite superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization.
+To this end, we model the detailed current density distribution inside the strand with a FE method, the CATI method [30] implemented in FiQuS [31, 32, 33], and we analyze the different dynamics resulting from the strand composite structure.
 2.1 Problem definition
-We consider a strand of diameter $d$ represented in Fig.
+We consider a multifilamentary strand such as in Fig.
 1.
-It consists of $N_{\text{f}}$ Nb-Ti filaments twisted with twist pitch length $p$ and embedded in a conducting copper (Cu) matrix.
-The geometrical parameters of the strand are summarized in Table 1.
-The strand carries no transport current but is subject to an applied (app) transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of amplitude $h_{\text{max}}$ (A/m), frequency $f$ (Hz), and direction $\hat{\boldsymbol{e}}_{y}$.
-Figure 1: Composite strand with Nb-Ti filaments embedded and twisted in a copper matrix (cylindrical outline) subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)$.
-Number of filaments ($N_{\text{f}}$) 54
-Filament diameter $90~{}\upmu$m
-Filament center-to-center spacing $110~{}\upmu$m
-Strand diameter ($d$) $1$ mm
-Twist pitch length ($p$) $19$ mm
-Table 1: Geometrical parameters of the strand.
-We fix the temperature to $1.9$ K.
-The resistivity of the Nb-Ti filaments is described by the power law [9] with power index $n=30$ and the field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$ (A/m2) given in Fig.
-2, with $\boldsymbol{b}$ (T) the local magnetic flux density, defined as $\boldsymbol{b}=\mu_{0}\boldsymbol{h}$, with $\mu_{0}=4\pi\times 10^{-7}$ H/m the permeability of vacuum.
-The resistivity of the copper matrix, accounting for magneto-resistance, is obtained from the STEAM material library [30], with a residual resistivity ratio $\text{RRR}=100$.
-Figure 2: Field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$. The maximum value is $j_{\text{c}}(\boldsymbol{0})=2.4\times 10^{10}$ A/m2.
-We model the strand response with the CATI method, proposed in [31].
-This method accounts for 3D effects due to the strand twist by using a pair of 2D models solved on a cross section of the strand, and coupled via circuit equations.
-As it is based on 2D models, it offers fast simulations.
-The method is implemented in GetDP [32] within FiQuS [33].
-FiQuS is developed at CERN as part of the STEAM framework [34].
-Geometry and mesh are performed by Gmsh [35].
-All the software is open-source and free to use.
-The code is available online [^2] https://gitlab.cern.ch/steam/analyses/cati-strand.
-2.2 Magnetization and loss
-We focus on magnetization and power loss. These are the two quantities we want to reproduce directly with the hysteresis model without having to compute the detailed current density distribution.
+It consists of twisted superconducting filaments embedded in a cylindrical conducting matrix.
+The strand carries no transport current but is subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$ of amplitude $h_{\text{max}}$ (A/m) and frequency $f$ (Hz).
+Figure 1: Composite strand with twisted superconducting filaments embedded in a conducting matrix (cylindrical outline) subject to a time-varying transverse magnetic field $\boldsymbol{h}_{\text{app}}(t)$.
+2.2 Magnetization and power loss
+We focus on magnetization and power loss. These are the two quantities to be reproduced with the ROHM model.
 The average magnetization vector $\boldsymbol{m}$ (A/m) is defined as the magnetic dipole moment per unit volume [34]. Based on the current density distribution on a given cross section, we have
 $\boldsymbol{m}=2\cdot\frac{1}{2a}\int_{S}\boldsymbol{x}\times\boldsymbol{j}\,\text{d}S,$ (1)
-with $a=\pi d^{2}/4$ the surface area of the strand cross section $S$, $\boldsymbol{x}$ the position vector, and $\boldsymbol{j}$ the current density.
-The multiplication by $2$ accounts for current loops closing at infinity in this 2D problem, as justified in A.
-The instantaneous power loss per unit length $P_{\text{tot}}$ (in W/m) is obtained by integrating the instantaneous power loss density, or Joule loss, over the strand cross section, including the Nb-Ti filaments and the Cu matrix,
+with $a$ the surface area of the strand cross section $S$, $\boldsymbol{x}$ the position vector, and $\boldsymbol{j}$ the current density.
+The multiplication by $2$ accounts for current loops closing at infinity, as justified in A.
+The instantaneous power loss per unit length $P_{\text{tot}}$ (W/m) is obtained by integrating the instantaneous power loss density, or Joule loss density, over the strand cross section, including the filaments and the matrix,
 $P_{\text{tot}}=\int_{S}\boldsymbol{j}\cdot\boldsymbol{e}\,\text{d}S,$ (2)
-with $\boldsymbol{e}=\rho\boldsymbol{j}$ the electric field (and $\rho$ the resistivity). The power loss per cycle and per unit length (in J/m) is evaluated as
-[…]
-To make the interpretation of the loss easier, and also because it will be helpful for the construction of the hysteresis model, the total power loss per cycle $Q_{\text{tot}}$ is decomposed in three contributions, as was done in [31]: hysteresis, coupling, and eddy loss.
+with $\boldsymbol{e}=\rho\boldsymbol{j}$ the electric field and $\rho$ the resistivity. The energy loss per cycle and per unit length (J/m) is evaluated as
+$Q_{\text{tot}}=\int_{1/f}^{2/f}P_{\text{tot}}(t)\,\text{d}t.$ (3)
+To make the interpretation of the loss easier, and to aid the construction of the ROHM model, the total energy loss per cycle $Q_{\text{tot}}$ is decomposed in three contributions: hysteresis, coupling, and eddy loss.
 The hysteresis loss is the loss induced by currents flowing in the superconducting filaments.
-The coupling loss is the loss induced by currents flowing in the conducting matrix, perpendicular to $\hat{\boldsymbol{e}}_{z}$ (the coupling currents).
+The coupling loss is the loss induced by currents flowing in the conducting matrix perpendicular to $\hat{\boldsymbol{e}}_{z}$ (the coupling currents).
 Finally, the eddy loss is the loss induced by currents flowing in the conducting matrix along $\hat{\boldsymbol{e}}_{z}$.
-Magnetization and loss are closely related to each other.
+Magnetization and loss are related to each other.
 Indeed, the area inside a closed magnetization loop is proportional to the total loss per cycle [35, 36].
@@
 2.3 Dynamic response of the strand
-The strand response is computed by the CATI method for a range of field amplitudes $h_{\text{max}}$ and frequencies $f$.
-The power loss per cycle induced by the transverse field is represented in Fig.
-3 as a function of $f$ for three different values of $h_{\text{max}}$.
-The $y$-component of the magnetization vector, $m$, is represented in Figs.
-4 and 5 for $\mu_{0}h_{\text{max}}=2$ T as a function of $\mu_{0}h_{\text{app}}(t)$ for selected frequencies.
+We now describe the strand response as a function of the field amplitude and frequency, and highlight its hysteretic behavior as well as the associated rate-dependency.
+For illustration, we consider a strand made of $54$ Nb-Ti filaments (diameter $90$ $\upmu$m) twisted with a twist pitch length of $19$ mm and embedded in a Cu matrix with a diameter of $1$ mm (same geometry as in [30]).
+We fix the temperature to $1.9$ K.
+The resistivity of the Nb-Ti filaments is described by the power law [9] with power index $n=30$ and a field-dependent critical current density $j_{\text{c}}(\boldsymbol{b})$ [37], with $\boldsymbol{b}$ (T) the local magnetic flux density, defined as $\boldsymbol{b}=\mu_{0}\boldsymbol{h}$, with $\mu_{0}=4\pi\times 10^{-7}$ H/m.
+The resistivity of the copper matrix accounts for magneto-resistance, with a residual resistivity ratio $\text{RRR}=100$ at zero field [37].
+2.3.1 Rate-independent regime
+At low frequencies (here for $f\lesssim 0.01$ Hz), the rate of change of the transverse magnetic field generates negligible coupling and eddy currents in the matrix.
+In this regime, the superconducting filaments behave almost uncoupled and most of the loss is due to hysteresis in them.
+A typical magnetization loop is shown in Fig.
+2.
+The magnetization decreases when the field increases as a result of the field-dependent critical current density.
+Strictly speaking, the finite value of $n$ creates a small rate-dependency, but in good approximation, the response can be considered rate-independent.
+A rate-independent ROHM model can therefore be used to describe the strand response in this regime.
+Figure 2: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T, $f=0.01$ Hz.
+2.3.2 Rate-dependent regime
+When the frequency increases, the response of the strand is no longer rate-independent.
+As shown in Fig.
+3, coupling and eddy loss can no longer be neglected, and the hysteresis loss varies significantly with frequency.
+As a result, the total loss per cycle is strongly rate-dependent.
+The magnetization also strongly depends on the frequency, as shown in Fig.
+4, where the $y$-component of the magnetization vector, $m$, is shown as a function of $\mu_{0}h_{\text{app}}(t)$ for different frequencies.
+To reproduce magnetization and loss in this regime, there is a clear need for a rate-dependent ROHM model.
+The design of such a model benefits from a good understanding of the physics behind the curves in Figs.
+3 and 4, which we briefly summarize below.
 Figure 3: Total loss per cycle and its contributions as a function of the frequency $f$, for three different magnetic field amplitudes.
@@
 Results from [30].
-At low frequencies, when the rate of change of the transverse magnetic field is not high enough for significant coupling currents and eddy currents to take place in the matrix, the superconducting filaments behave close to uncoupled.
-In that regime, most loss is due to hysteresis in the filaments.
-A typical magnetization loop in that case is represented in Fig.
-4.
-The magnetization decreases when the field increases as a result of the field-dependent critical current density.
-In this regime, a good first approximation is to consider that the response is rate-independent (it is not strictly the case in reality due to the finite value of the $n$ index in the power law).
-To this end, a rate-independent hysteresis model can be used.
-Figure 4: Strand magnetization along $\hat{\boldsymbol{e}}_{y}$ for $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. Same result as the first curve in Fig. 5.
-When the frequency increases, the response of the strand becomes strongly rate-dependent.
-As shown in Fig.
-3, coupling and eddy loss can dominate the total loss, and the filament loss is not constant with frequency.
-As a result, the loss per cycle can exhibit more than one order of magnitude of variation as a function of frequency for an identical applied field amplitude.
-Similarly, the magnetization strongly depends on the frequency, as shown in Fig.
-5.
-To reproduce these regimes, there is a clear need for a rate-dependent hysteresis model.
-The design of such a model benefits from a good understanding of the physics of the curves in Figs.
-3 and 5.
-Figure 5: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T and for different frequencies $f$. Results from [31] using Eq. (1).
+Figure 4: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T and for different frequencies $f$. Results from [30] using Eq. (1). The curve for $f=0.01$ Hz is shown in more detail in Fig. 2.
 For low field amplitudes, the dynamics of the coupling currents and of the associated loss is well described by analytical models [5, 38].
-The coupling loss follows follows a typical bell curve with a maximum at a characteristic frequency $f_{\text{c}}$.
+The coupling loss follows a typical bell curve with a maximum at a frequency $f_{\text{c}}$.
 With increasing field amplitude, the peak of the coupling loss curve shifts towards lower frequencies, as can be seen in Fig.
 3 for $\mu_{0}h_{\text{max}}=2$ T.
-This is associated with the saturation of superconducting filaments, which limit the coupling currents amplitude, and to a change of the effective permeability, due to the loss of diamagnetic effect due to saturated filaments, as discussed in [5].
-The coupling currents also give rise to a change of regime for the filament magnetization and loss.
-As mentioned previously, at low frequencies, $f\lesssim f_{\text{c}}$, they are mostly uncoupled, producing a relatively small magnetization.
-On the other hand, for $f\gtrsim f_{\text{c}}$, they are coupled and exchange currents via coupling currents.
-As a result, their magnetization is increased, see for example the situation for $f=10$ Hz in Fig.
-5.
-Whether the filament loss for coupled filaments is higher or lower than that for uncoupled filaments depends on $\mu_{0}h_{\text{max}}$, as can be seen in Fig.
+This is associated with the saturation of superconducting filaments, which limits the coupling currents amplitude, and to the loss of the diamagnetic effect in saturated filaments, as discussed in [5].
+The coupling current dynamics influences the filament magnetization and hysteresis loss.
+At low frequencies, $f\lesssim f_{\text{c}}$, filaments are mostly uncoupled, whereas at higher frequencies, $f\gtrsim f_{\text{c}}$, they exchange currents via coupling currents.
+As a result, the overall magnetization increases, see for example the situation for $f=10$ Hz in Fig.
+4, and the hysteresis loss is affected.
+Whether the hysteresis loss per cycle for coupled filaments is higher or lower than that for uncoupled filaments depends on the applied field amplitude, as can be seen in Fig.
 3.
 At even higher frequencies, eddy currents in the conducting matrix become the dominating factor for loss and magnetization.
-As a result of eddy currents and the skin effect, the inner part of the strand, containing the filaments, is shielded from the outer field and the filament loss decreases accordingly.
-In that regime, the magnetization curve approaches the shape of an ellipse [37] that gets thinner and thinner with increasing frequency, and of slope approaching $-2$ because of demagnetization effects (this is further explained in Section 3.1[38].
+As a result of eddy currents and the skin effect, the inner part of the strand, containing the filaments, is shielded from the external field and the hysteresis loss decreases accordingly.
+In that regime, the magnetization curve approaches the shape of an ellipse [35] that becomes thinner and thinner with increasing frequency, and its slope approaches $-2$ because of demagnetization effects [36].
 3 Reduced Order Hysteretic Magnetization Model
-The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method.
-This is fast and convenient enough to get a good understanding of a single strand response.
-However, for simulating a full-scale magnet, containing thousands of strands, one cannot afford such a detailed description of the current density at the strand level.
-Instead, one has to use an alternative model that produces similar macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but without requiring a detailed calculation of the current density, so as to strongly reduce the computational cost.
-This idea is illustrated in Fig.
-6.
-We present such a model in this section, the Reduced Order Hysteretic Magnetization (ROHM) model.
-[…]
-This model is adapted and inspired from the energy-based model for ferromagnetic materials [27, 28].
+The simulation time for each of the simulations presented in the previous section is of the order of one hour with the CATI method [30].
+This is fast and convenient enough to get a good understanding of a single strand response with a detailed description of the current density distribution (as discussed in [30] or shown later in Fig.
+27).
+However, to simulate transients in a full-scale magnet containing thousands of strands, one cannot afford models involving such a detailed description at the strand level.
+Instead, one can use an alternative model that produces accurate macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but that does not rely on a detailed calculation of the current density.
+We present such a model in this section: the Reduced Order Hysteretic Magnetization (ROHM) model.
+This model is adapted from the energy-based model for ferromagnetic materials [27, 28].
 We start in Section 3.1 by clarifying how the ROHM model can be used to model the strand.
-We then define the building blocks, referred to as cells, that will be combined for the construction of the model.
+We then define the building blocks, referred to as cells, that will be combined for the construction of the ROHM model.
 Each cell defines a local hysteretic relationship between the magnetic field $\boldsymbol{h}$ and the magnetic flux density $\boldsymbol{b}$.
@@
 In Section 3.3, we add contributions to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell).
-In Section 3.4, we discuss how to reproduce the effect of a field-dependent critical current density within the cells.
-Finally, we show in Section 3.5 how several cells can be combined into a chains of cell, defining a complete ROHM model.
-3.1 Concept of a reduced order model
+Finally, we show in Section 3.4 how to combine several cells into a chain of cells, defining a complete ROHM model.
+3.1 Homogenized model and internal field
 […]
+Figure 5: Homogenization concept without transport current.
 […]
 $\boldsymbol{b}=\mu_{0}\boldsymbol{h}+\mu_{0}\boldsymbol{m}.$ (5)
-Similarly, the power loss is no longer obtained via a current density as in Eq.
-(2), but is now instead described by the magnetic power density $\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$.
-Part of this power is related to stored magnetic energy, and hence not associated with loss, and the remaining part is related to dissipated energy.
-One major advantage of the energy-based hysteresis model used in this work is that it gives access to both parts separately at any time [27].
+Similarly, the power loss is no longer computed as a Joule loss as in Eq.
+(2), but is now linked with the magnetic power density $p_{\text{tot}}=\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$.
+Part of this power is related to stored magnetic energy, and hence is not associated with loss.
+The complementary part is related to dissipated energy.
+One major advantage of the energy-based hysteresis model used in this work is that it gives access to both parts (stored and dissipated) separately at all time [27].
 […]
@@
 $\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\frac{1}{2}\boldsymbol{m},$ (7)
-where $\boldsymbol{h}_{\text{in}}$ has no immediate meaning in the detailed strand model per se, but is the local magnetic field the hysteresis model will see inside the strand.
+where $\boldsymbol{h}_{\text{in}}$ has no immediate meaning in the detailed strand model per se, but is the local magnetic field that the ROHM model will see inside the strand.
 These curves are shown in Fig.
-7.
+6.
 At high frequencies ($f\gtrsim 10^{4}$ Hz), they approach the shape of an ellipse with a slope of $-1$.
-The area of closed loops are identical in both Figs.
-5 and 7, as shown in B.
-Figure 7: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T. Same results as in Fig. 5 but in terms of $\mu_{0}h_{\text{in}}$ instead of $\mu_{0}h_{\text{app}}$ (see Eq. (7)).
+The area of the closed loops is preserved between Fig.
+4 and Fig.
+6, as shown in B.
+Figure 6: Strand magnetization for $\mu_{0}h_{\text{max}}=2$ T. The results are the same as those in Fig. 4 but in terms of $\mu_{0}h_{\text{in}}$ instead of $\mu_{0}h_{\text{app}}$ (see Eq. (7)).
 3.2 Superconductor cell - S cell
-As a first step, we present a rate-independent hysteresis cell [27], that we refer to as the superconductor cell, or S cell.
-In this cell, the magnetic field $\boldsymbol{h}$ is decomposed into a reversible field, $\boldsymbol{h}_{\text{rev}}$, and an irreversible field, $\boldsymbol{h}_{\text{irr}}$:
+As a first step, we present a rate-independent hysteresis cell [27]: the superconductor cell, or S cell.
+In this cell, the magnetic field $\boldsymbol{h}$ is decomposed into a reversible field, $\boldsymbol{h}_{\text{rev}}$, and an irreversible field [^2] The irreversible field $\boldsymbol{h}_{\text{irr}}$ has nothing to do with the irreversibility field of a superconductor., $\boldsymbol{h}_{\text{irr}}$:
 $\boldsymbol{h}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{irr}}.$ (8)
-The reversible field defines the magnetic flux density:
+The reversible field defines the magnetic flux density
 $\boldsymbol{b}=\mu_{0}\boldsymbol{h}_{\text{rev}}.$ (9)
@@
 Its amplitude $\|\boldsymbol{h}_{\text{irr}}\|$ is bounded by a value $\kappa$ (A/m), the irreversibility parameter.
-From a known magnetic field $\boldsymbol{h}$, which is the driving vector, and a given reversible field $\boldsymbol{h}_{\text{rev}}$ that has been established previously (as a function of the history of $\boldsymbol{h}$), the irreversible field $\boldsymbol{h}_{\text{irr}}$ is determined as follows:
+From a known magnetic field $\boldsymbol{h}$, and a given reversible field $\boldsymbol{h}_{\text{rev}}$ that has been established previously (as a function of the history of $\boldsymbol{h}$), the irreversible field $\boldsymbol{h}_{\text{irr}}$ is determined as follows:
 $\boldsymbol{h}_{\text{irr}}$ =$&\boldsymbol{h}-\boldsymbol{h}_{\text{rev}},\quad&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|<\kappa,\\ &\kappa\ \dot{\boldsymbol{b}}/\|\dot{\boldsymbol{b}}\|,\quad&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|=\kappa,\\ \end{aligned}\right.$ (10)
-where the dot notation (as in $\dot{\boldsymbol{b}}$) represents the time derivative of the quantity.
+where the dot notation (as in $\dot{\boldsymbol{b}}$) represents the time derivative.
 An illustration of this equation is given in Fig.
-8.
-While the driving field $\boldsymbol{h}$ evolves inside the sphere of radius $\kappa$ centered at $\boldsymbol{h}_{\text{rev}}$, the reversible field $\boldsymbol{h}_{\text{rev}}$ stays constant, and the irreversible field is defined accordingly, see Fig.
-8(a).
-If $\boldsymbol{h}$ reaches the surface of the sphere, then, the reversible field $\boldsymbol{h}_{\text{rev}}$ must evolve as well, in order to maintain the condition $\|\boldsymbol{h}_{\text{irr}}\|\leq\kappa$ valid, and $\boldsymbol{h}_{\text{irr}}$ is established parallel to $\dot{\boldsymbol{b}}$, see Fig.
-8(b).
+7.
+While the driving field $\boldsymbol{h}$ evolves inside the sphere of radius $\kappa$ centered at $\boldsymbol{h}_{\text{rev}}$, the reversible field $\boldsymbol{h}_{\text{rev}}$ stays constant (and so does $\boldsymbol{b}=\mu_{0}\boldsymbol{h}_{\text{rev}}$), and the irreversible field is defined accordingly, see Fig.
+7(a).
+If $\boldsymbol{h}$ reaches the surface of the sphere, then, the reversible field $\boldsymbol{h}_{\text{rev}}$ (and, hence, also $\boldsymbol{b}$) must evolve in order to maintain the condition $\|\boldsymbol{h}_{\text{irr}}\|\leq\kappa$ valid, and $\boldsymbol{h}_{\text{irr}}$ is established parallel to $\dot{\boldsymbol{b}}$, see Fig.
+7(b).
 (a) $\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev}}\|<\kappa$.
@@
 Figure 7: S cell, adapted from [27, 29].
-Illustration of the magnetic field decomposition into reversible and irreversible fields in 2D.
+Illustration in 2D of the magnetic field decomposition into reversible and irreversible fields.
 The dotted elements in (b) refer to the situation at a previous time.
-The total power density $p_{\text{tot}}$ (in W/m3) is expressed as
+The total power density $p_{\text{tot}}$ (W/m3) is expressed as
 $p_{\text{tot}}$ $=\boldsymbol{h}\cdot\dot{\boldsymbol{b}}$
 $=\underbrace{\boldsymbol{h}_{\text{rev}}\cdot\dot{\boldsymbol{b}}}_{p_{\text{rev}}}+\underbrace{\boldsymbol{h}_{\text{irr}}\cdot\dot{\boldsymbol{b}}}_{p_{\text{irr}}}.$ (11)
-The reversible power $p_{\text{rev}}$ is the time derivative of a stored magnetic energy density,
+The reversible power $p_{\text{rev}}$ is the time derivative of the stored magnetic energy density,
 $p_{\text{rev}}=\boldsymbol{h}_{\text{rev}}\cdot\mu_{0}\dot{\boldsymbol{h}}_{\text{rev}}=\frac{d}{dt}\left(\frac{1}{2}\boldsymbol{h}_{\text{rev}}\cdot\mu_{0}\boldsymbol{h}_{\text{rev}}\right),$ (12)
-while the irreversible power $p_{\text{irr}}$ is always non-negative and is associated with dissipated energy density,
+while the irreversible power $p_{\text{irr}}$ is always non-negative and is associated with the dissipated energy density,
 $p_{\text{irr}}=\boldsymbol{h}_{\text{irr}}\cdot\dot{\boldsymbol{b}}=\kappa\|\dot{\boldsymbol{b}}\|=\kappa\|\mu_{0}\dot{\boldsymbol{h}}_{\text{rev}}\|.$ (13)
 It is one of the main benefits of the energy-based approach that the dissipated energy is clearly separated from the stored energy [27].
-A mechanical analogy of a S cell in a one-dimensional setting is shown in Fig.
-9.
-It represents a mechanical system made of a dry friction element in parallel with a restoring spring [27].
-Examples of hysteresis curves computed with the S cell are shown in Fig.
-10.
-These curves are independent of the rate of change of the applied field.
+A mechanical analogy of the S cell in a one-dimensional setting is shown in Fig. 8. It represents a mechanical system made of a dry friction element in parallel with a restoring spring [27].
 Figure 8: Mechanical analogy [27] of the S cell: a parallel connection of a dry friction element and a restoring spring.
-The system is subject to a force $h$, which causes an elongation $b=\mu_{0}h_{\text{rev}}$.
-Figure 10: Hysteresis curves for the S cell subject to a unidirectional oscillating magnetic field $h$ as an input.
+In the analogy, the system is subject to a force $h$ which causes an elongation $b=\mu_{0}h_{\text{rev}}$.
+In practice, the magnetization of superconductors decreases with increasing field as a consequence of the decreasing critical current density $j_{\text{c}}$, as is shown in Fig.
+2.
+In order to reproduce this effect with the S cell, one can make the irreversibility parameter $\kappa$ a function of the magnetic flux density $\boldsymbol{b}$, by defining
+$\kappa(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa},$ (14)
+with $\bar{\kappa}$ (A/m) a constant and $f_{\kappa}$ a scaling function, smoothly evolving from $1$ at zero field to 00 at large fields.
+The expression of $f_{\kappa}$ is function of the geometry and must be determined at the parameter identification step, as discussed in Section 5.
+The introduction of a field-dependent irreversibility parameter brings an additional nonlinearity in the equations, which has to be addressed during the numerical simulation.
+Magnetization loops obtained from S cells with different irreversibility parameters are shown Fig.
+9 for a unidirectional excitation.
+By construction, these curves are independent of the rate of change of the applied field.
+Figure 9: Hysteresis curves obtained with one S cell with field-dependent or field-independent irreversibility parameters (the function $f_{\kappa}(b)$ is the same as in Section 6).
 The legend is the same for both subfigures.
-(Left) Magnetic flux density $b$ versus applied field $\mu_{0}h$.
-(Right) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
+(Up) Magnetic flux density $b$ versus applied field $\mu_{0}h$.
+(Down) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
+Note that in this simple case, $h_{\text{irr}}=-m$.
+A careful choice of the scaling function $f_{\kappa}(b)$ allows to reproduce the shape of the lower and upper branches of the major hysteresis loop of Fig.
+2.
+However, with a single S cell, the transitions between these upper and lower branches are not smooth as in the reference solution in Fig.
+2, but sharp.
+Also, for applied fields smaller than $\kappa(0)$, the magnetic flux density remains exactly zero and no loss is generated by the S cell.
+To obtain smooth transitions and better description at small field amplitudes, combining several cells is necessary.
+This is discussed in Section 3.4.
 3.3 Composite superconductor cell - CS cell
-In regimes in which the superconducting strand response is clearly not rate-independent, the S cell presented in the previous section must be generalized.
+For regimes in which the superconducting strand response is strongly rate-independent, the S cell presented in the previous section must be generalized.
 Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq.
@@
 The magnetic field decomposition is represented in Fig.
-11(b).
+10(b).
 As described below, the fields $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$ are always parallel to $\dot{\boldsymbol{b}}=\mu_{0}\dot{\boldsymbol{h}}_{\text{rev}}$.
 We refer to this new cell as the composite superconductor cell, or CS cell.
-Adapting [28], the field $\boldsymbol{h}_{\text{eddy}}$ is defined as
+(a) Mechanical analogy.
+(b) Illustration in 2D.
+Figure 10: CS cell generalizing the S cell by accounting for eddy currents and coupling current effects, as well as for filament coupling effect.
+Illustration of the magnetic field decomposition into four contributions ($\boldsymbol{h}_{\text{rev}}$, $\boldsymbol{h}_{\text{irr}}$, $\boldsymbol{h}_{\text{eddy}}$, and $\boldsymbol{h}_{\text{coupling}}$).
+The dashed circle of radius $\chi$ limits the amplitude of the field $\boldsymbol{h}_{\text{coupling}}$.
+The dotted arrow represents $\boldsymbol{h}_{\text{rev}}$ at a previous time.
+Adapting [28], the field $\boldsymbol{h}_{\text{eddy}}$ is defined by
 $\boldsymbol{h}_{\text{eddy}}=\tau_{\text{e}}\dot{\boldsymbol{h}}_{\text{rev}},$ (17)
-with an eddy current time constant parameter $\tau_{\text{e}}$ (s). This component produces a loss contribution, $p_{\text{eddy}}$, which is non-negative. It reads:
+with an eddy current time constant parameter $\tau_{\text{e}}$ (s). The eddy component produces a new loss contribution, $p_{\text{eddy}}$ (W/m3), which reads
 $p_{\text{eddy}}=\boldsymbol{h}_{\text{eddy}}\cdot\dot{\boldsymbol{b}}=\frac{\mu_{0}}{\tau_{\text{e}}}\|\boldsymbol{h}_{\text{eddy}}\|^{2}=\frac{\tau_{\text{e}}}{\mu_{0}}\|\dot{\boldsymbol{b}}\|^{2}.$ (18)
 In the mechanical analogy, the inclusion of eddy current effects corresponds to adding a dashpot (viscous friction element) in parallel with the dry friction element and the restoring spring, as illustrated in Fig.
-11(a).
-For slow variations of the force $h$, the dashpot is barely opposing the elongation $b$, whereas for faster variations of $h$, it becomes gradually more blocking and starts to oppose to variations of $b$.
+10(a).
+For slow variations of the force $h$, the dashpot is barely opposing the elongation $b$, whereas for faster variations of $h$, it becomes gradually more blocking and starts to oppose variations of $b$.
 This reproduces magnetic shielding and increased magnetization due to eddy currents.
-The next new contribution $\boldsymbol{h}_{\text{coupling}}$ is associated with coupling currents [5].
-The coupling currents in the conducting matrix are fed by currents flowing in the superconducting filaments, and are therefore also subject to hysteresis and saturation of the filaments.
-The field $\boldsymbol{h}_{\text{coupling}}$ has then its own irreversibility parameter $\chi$ (in A/m).
+The other new contribution $\boldsymbol{h}_{\text{coupling}}$ is associated with coupling currents [5].
+The coupling currents in the conducting matrix are due to net currents flowing in the superconducting filaments, and are therefore also subject to hysteresis and saturation of the filaments.
+The field $\boldsymbol{h}_{\text{coupling}}$ has then its own irreversibility parameter $\chi$ (A/m).
 We define
-$\boldsymbol{h}_{\text{coupling}}$ =$&\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}},&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|<\chi,\\ &\chi\ \dot{\boldsymbol{h}}_{\text{rev}}/\|\dot{\boldsymbol{h}}_{\text{rev}}\|,&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\geq\chi.\\ \end{aligned}\right.$ (18)
-$=\frac{\tau_{\text{c}}}{\mu_{0}}\dot{\boldsymbol{b}}_{\text{coupling}},$ (19)
-with a coupling currents time constant $\tau_{\text{c}}$ (s).
-The field $\boldsymbol{h}_{\text{coupling}}$ introduces two new loss contributions, $p_{\text{coupling}}$ and $p_{\text{irr,c}}$, that read:
+$\boldsymbol{h}_{\text{coupling}}$ =$&\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}},&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|<\chi,\\ &\chi\ \dot{\boldsymbol{h}}_{\text{rev}}/\|\dot{\boldsymbol{h}}_{\text{rev}}\|,&&\text{if }\|\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}\|\geq\chi,\\ \end{aligned}\right.$ (19)
+with a coupling current time constant $\tau_{\text{c}}$ (s).
+In order to reproduce reduced magnetization at higher fields, the irreversibility parameter $\chi$ must also be a function of $\boldsymbol{b}$.
+We therefore define
+$\chi(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi},$ (20)
+with $\bar{\chi}$ a constant (A/m) and $f_{\chi}(\boldsymbol{b})$ a scaling function.
+As for $f_{\kappa}(\boldsymbol{b})$, the expression of $f_{\chi}(\boldsymbol{b})$ must be determined in the parameter identification step.
+The field $\boldsymbol{h}_{\text{coupling}}$ introduces two new loss contributions, $p_{\text{coupling}}$ and $p_{\text{irr,c}}$ (W/m3), that read
 $p_{\text{coupling}}$ $=\frac{\mu_{0}}{\tau_{\text{c}}}\|\boldsymbol{h}_{\text{coupling}}\|^{2}=\frac{\tau_{\text{c}}}{\mu_{0}}\|\dot{\boldsymbol{b}}_{\text{coupling}}\|^{2},$ (21)
-$p_{\text{irr,c}}$ $=\boldsymbol{h}_{\text{coupling}}\cdot\left(\dot{\boldsymbol{b}}-\dot{\boldsymbol{b}}_{\text{coupling}}\right).$ (21)
-The mechanical analogy associated with this new problem is represented in Fig.
-11(a).
+$p_{\text{irr,c}}$ $=\boldsymbol{h}_{\text{coupling}}\cdot\left(\dot{\boldsymbol{b}}-\dot{\boldsymbol{b}}_{\text{coupling}}\right),$ (22)
+with
+$\dot{\boldsymbol{b}}_{\text{coupling}}=\frac{\mu_{0}}{\tau_{\text{c}}}\boldsymbol{h}_{\text{coupling}}.$ (23)
+The mechanical analogy associated with the coupling current contribution is represented in Fig.
+10(a).
 The combination of the dashpot in series with the dry friction element reproduces the saturation of the coupling currents.
@@
 If the force exceeds $\chi$ on the dashpot, it is then subject to hysteresis.
-The second hysteresis element, with $\chi$, always acts in parallel with the first one, related to filament magnetization.
-The magnetization of coupled filaments therefore depends of the sum of both.
-(a) Mechanical analogy.
-(b) Illustration in 2D.
-Figure 11: CS cell accounting for eddy currents and coupling current effects. The dashed circle of radius $\chi$ limits the amplitude of the field $\boldsymbol{h}_{\text{coupling}}$.
+The second hysteresis element, with $\chi$, always acts in parallel with the first one, related to the irreversible field $\boldsymbol{h}_{\text{irr}}$.
+The magnetization of coupled filaments therefore depends on the sum of both, and so is the associated hysteresis loss.
+We therefore group $p_{\text{irr}}$ and $p_{\text{irr,c}}$ into a single component $p_{\text{hyst}}=p_{\text{irr}}+p_{\text{irr,c}}$.
 Note that by choosing $\tau_{\text{e}}=0$ and $\tau_{\text{c}}=0$, we retrieve the S cell of Section 3.2. The S cell is therefore a particular case of the CS cell.
-An illustration of magnetization curves obtained with a CS cell is given in Fig.
-12, and the loss per cycle as a function of frequency is given in Fig.
-13.
-The loss per cycle curve exhibits the sought features for the composite strand.
-The filament curve shows the expected evolution, two plateaus at low and middle frequencies, with a transition where the coupling loss peaks, and then a sharp decrease when the eddy loss becomes dominant.
-Coupling and eddy loss curves both follow the shape of two bell curves.
-Figure 12: Hysteresis curves for the composite hysteresis cell reproducing eddy and coupling currents, with $\kappa=0.2/\mu_{0}$ A/m, $\chi=0.3/\mu_{0}$ A/m, $\tau_{\text{c}}=0.1$ s, and $\tau_{\text{e}}=70$ $\upmu$s.
-Unidirectional oscillating magnetic field $h$ as an input.
-The legend is shared between both subfigures.
-(Left) Magnetic flux density $b$ versus applied field $\mu_{0}h$.
-(Right) Magnetization $\mu_{0}m=b-\mu_{0}h$ versus applied field $\mu_{0}h$.
-Figure 13: Total loss per cycle and separated loss contributions as a function of the frequency $f$, for a magnetic field amplitude of $2$ T obtained with the same composite hysteresis cell as in Fig.
-12.
-Note that the eddy current curve for the eddy loss per cycle should be asymptotically decreasing as $1/\sqrt{f}$ at high frequencies, due to the skin effect.
-With the CS cell, one can however only expect the curve to decrease as $1/f$ at high frequencies.
-For $f\to\infty$, the CS cell is therefore not expected to provide perfectly valid results.
-It should be further improved if a more realistic frequency response is wanted in such high frequencies.
-For most practical applications however, the field rates of change are sufficiently low for this error to stay small.
-Comparing with the results from a detailed multifilamentary strand model in Fig.
-3, one can see that a single CS cell is still not sufficiently accurate.
-Several of them have to be combined to achieve a better matching with the reference, as will be discussed in Section 3.5.
-3.4 Inclusion of $j_{\text{c}}(b)$ field dependence
-In reality, the magnetization of superconductors decreases with increasing field, as a consequence of the decreasing critical current density $j_{\text{c}}$.
-Such an effect is not yet reproduced with the hysteresis cells defined above, because the irreversibility parameters $\kappa$ and $\chi$ are still constant values.
-By making the irreversibility parameters $\kappa$ and $\chi$ decrease with increasing field, one can produce more realistic magnetization curves.
-In general, we therefore define field-dependent irreversibility parameters:
-$\kappa(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa},\quad\chi(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi},$ (22)
-with $\bar{\kappa}$ and $\bar{\chi}$ two constants (in A/m) and $f_{\kappa}$ and $f_{\chi}$ two scaling functions, smoothly evolving from $1$ at zero field to 00 at large fields.
-The expressions of $f_{\kappa}$ and $f_{\chi}$ are functions of the geometry and must be determined at the parameter identification step.
-This will be discussed in Section 5.
-Note that the introduction of field-dependent irreversibility parameters in the cells brings an additional nonlinearity to the equations to be addressed during the numerical simulation.
+In summary, the power loss decomposition reads:
+$p_{\text{tot}}=p_{\text{rev}}+\underbrace{p_{\text{irr}}+p_{\text{irr,c}}}_{p_{\text{hyst}}}+p_{\text{eddy}}+p_{\text{coupling}},$ (24)
+where $p_{\text{rev}}$ corresponds to stored energy, while the other components correspond to dissipated energy.
+Magnetization curves obtained with one CS cell and the associated energy loss per cycle and per unit volume are shown in Figs.
+11 and 12.
+For a given frequency $f$, the energy loss per cycle and per unit volume is computed as
+$q_{\text{tot}}=\int_{1/f}^{2/f}p_{\text{tot}}(t)\,\text{d}t.$ (25)
+The total loss $q_{\text{tot}}$ is also decomposed in three contributions for easier interpretation, following Eq. (24).
+The hysteresis loss exhibits the expected evolution with frequency: two plateaus at low and middle frequencies, with a transition where the coupling loss peaks, and then a sharp decrease when the eddy loss becomes dominant.
+Coupling and eddy loss curves follow the shape of two bell curves.
+Figure 11: Hysteresis curves obtained with one CS cell at different frequencies.
+Cell parameters are: $\mu_{0}\bar{\kappa}=0.2$ T, $\mu_{0}\bar{\chi}=0.3$ T, $\tau_{\text{c}}=0.2$ s, and $\tau_{\text{e}}=0.3$ ms, and the same functions $f_{\kappa}(b)$ and $f_{\chi}(b)$ as in Section 7.
+[…]
+In reality, the eddy current loss is asymptotically decreasing as $1/\sqrt{f}$ at high frequencies due to the skin effect.
+With the CS cell, one can however only expect the curve to decrease as $1/f$ at high frequencies, as a result of Eq.
+(17).
+For $f\to\infty$, the CS cell is therefore not expected to provide perfect results and shall be adapted if one wants better results for very high frequencies ($f\gtrsim 10^{3}$ Hz).
+For the same reasons as for the S cell, several CS cells must be combined to obtain a good description on a range of field amplitudes. This is discussed in the next section.
 3.4 Chain of cells
-The S cell and the CS cell contain the necessary ingredients to reproduce the different regimes observed in the multifilamentary strand, but a single cell alone does not provides a faithful description over a wide field range.
-Indeed, if the applied magnetic field is smaller than $\kappa$, both the S cell and the CS cell produce exactly zero magnetic flux density, and hence no loss.
-In reality, the evolution towards saturation is smooth and progressive, and not subject to a single threshold $\kappa$, as in Figs.
-10 and 12.
-To better approach the smooth magnetization curves of Figs. 4 and 5, we can combine several cells into what we call a chain of cells. To this end, we follow the approach proposed in [27].
-The approach consists of decomposing the total magnetic flux density $\boldsymbol{b}$ into a number of $N\in\mathbb{N}_{0}$ contributions $\alpha_{k}\boldsymbol{b}_{k}$, with $k=1,\dots,N$.
-The $\boldsymbol{b}_{k}$ are described by distinct hysteresis cells, with distinct parameter values and the $\alpha_{k}$ are weights associated with each of them.
-All cells are subject to the same total magnetic field $\boldsymbol{h}$ and contribute to the total magnetic flux density with weight $\alpha_{k}$.
+The S cell and the CS cell contain the necessary ingredients to reproduce the different regimes observed in a multifilamentary strand, but a single cell alone does not produce a valid response over a wide field range.
+Indeed, in reality, the evolution towards saturation is smooth and progressive, and not subject to a single threshold field $\kappa$ as is the case in Figs.
+9 and 11.
+To better approach the smooth magnetization curves of Figs. 2 and 4, we can combine several cells into a chain of cells. To this end, we follow the approach proposed in [27].
+The approach consists in decomposing the total magnetic flux density $\boldsymbol{b}$ into $N$ contributions $\alpha_{k}\boldsymbol{b}_{k}$, with $k=1,\dots,N$.
+Each $\boldsymbol{b}_{k}$ is described by a distinct hysteresis cell, but all cells are driven by the same magnetic field $\boldsymbol{h}$.
+Each cell contributes to the total magnetic flux density with a weight $\alpha_{k}$.
 We define
@@
 where $\boldsymbol{h}_{\text{rev},k}$ is the reversible field associated with cell $k$.
-With the mechanical analogy, this corresponds to connecting cells as a chain, i.e., in series, as illustrated in Fig.
-14.
-The power loss is computed cell by cell, in terms of the fractions $\alpha_{k}\dot{\boldsymbol{b}}_{k}$ of the total rate of change $\dot{\boldsymbol{b}}$.
-Figure 14: Chain of cells.
-The total magnetic flux density is computed as $\boldsymbol{b}=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{b}_{k}$.
-Arbitrary combinations of cells can be put in series.
-The represented arrangement is relevant for a multifilamentary strand, as explained in Section 7.
+In the mechanical analogy, this corresponds to connecting cells as a chain, i.e., in series, as illustrated in Fig.
+13.
+Figure 13: (Up) Chain of S cells. (Down) Chain of CS cells.
+The power loss is computed by adding contributions of each individual cell, involving the fractions $\alpha_{k}\dot{\boldsymbol{b}}_{k}$ of the total rate of change $\dot{\boldsymbol{b}}$.
 The advantage of this approach is that each cell can be solved independently, as they are all subject to the same magnetic field $\boldsymbol{h}$.
-Fixed point iterations can be performed to account for field-dependent parameters $\kappa$ or $\chi$ that should depend on the total magnetic flux density $\boldsymbol{b}$.
-The equations for each cell are identical to what was described in Sections 3.2 and 3.3, but each cell only contributes partly to the total magnetic flux density $\boldsymbol{b}$, with a fraction $\alpha_{k}\boldsymbol{b}_{k}$.
-Once parameters are identified (see Section 5), the chain of cells defines the hysteresis model, i.e., the constitutive relationship $\boldsymbol{b}=\mathcal{B}(\boldsymbol{h})$ that reproduces the composite strand magnetization and loss.
-4 Implementation and Inclusion in FE Framework
-The hysteresis model output the magnetic flux density $\boldsymbol{b}(t)$ as a function of a given magnetic field $\boldsymbol{h}(t)$ (from which we can compute the magnetization $\boldsymbol{m}=\boldsymbol{b}/\mu_{0}-\boldsymbol{h}$ and the power loss).
-Both $\boldsymbol{b}(t)$ and $\boldsymbol{h}(t)$ are vector quantities varying in time.
-For a numerical simulation, time is discretized and the solutions are sought at successive time steps, based on the knowledge of the solution at previous time steps and on the new value of the magnetic field.
-In this section, we first propose update rules for the two types of cells defined in the previous section, the S cell and CS cell.
-We then discuss the case of a chain of cells and conclude by describing how to include the hysteresis model in a FE framework.
+Fixed point iterations can be performed to account for field-dependent parameters $\kappa$ or $\chi$ that depend on the total magnetic flux density $\boldsymbol{b}$.
+The equations for each cell are identical to those in Sections 3.2 and 3.3, with the only difference that each cell only contributes partly to the magnetic flux density $\boldsymbol{b}$, with a fraction $\alpha_{k}\boldsymbol{b}_{k}$.
+[…]
+4 Implementation and Inclusion in a FE Framework
+The ROHM model gives the magnetic flux density $\boldsymbol{b}(t)$ as a function of a given magnetic field $\boldsymbol{h}(t)$, from which we can compute the magnetization $\boldsymbol{m}(t)=\boldsymbol{b}(t)/\mu_{0}-\boldsymbol{h}(t)$ and the power loss.
+All are vector quantities varying in time.
+For a numerical simulation, time is discretized and the solutions are computed at successive time steps, based on the knowledge of the solution at previous time steps and on the new value of the magnetic field.
+In this section, we show how to solve the two types of cells defined in the previous section, the S cell and CS cell, as well as chains of cells.
+We then describe how to include the ROHM model in a FE framework.
 4.1 S cell
 If the irreversibility parameter $\kappa$ is constant, the S cell can be solved with an explicit update rule, as proposed in [27].
-Let $\boldsymbol{h}_{\text{rev(p)}}$ be the reversible field computed at the previous time step.
-For a magnetic field $\boldsymbol{h}$, the updated reversible field $\boldsymbol{h}_{\text{rev}}$ reads:
+Let $\boldsymbol{h}_{\text{rev(p)}}$ be the reversible field computed at the previous time step, as a result of the history of $\boldsymbol{h}$.
+For a magnetic field $\boldsymbol{h}$ at the new time step, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated as follows:
 $\boldsymbol{h}_{\text{rev}}$ =$&\boldsymbol{h}_{\text{rev(p)}},&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},&&\text{otherwise,}\end{aligned}\right.$
@@
 (12) and (13), respectively.
-If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ as proposed in Section 3.4 to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit.
-In such a case however, we observed that a good approximate solution is found easily in a few fixed point iterations, that is, by solving successively Eq.
+If the irreversibility parameter $\kappa$ is not constant but a function of $\boldsymbol{b}$ to account for field-dependent $j_{\text{c}}$, the update rule is no longer explicit.
+In such a case, an approximate solution can be computed easily with a few fixed point iterations, that is, by solving successively Eq.
 (4.1) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
-It is remarkable that the implementation of the S cell entirely consists in the simple test of Eq.
+The implementation of the S cell is very simple, as it entirely consists in the test of Eq.
 (4.1).
-It only requires the knowledge of the previous reversible field, which has to be stored as an internal variable in the numerical implementation.
+It only requires the knowledge of the previous reversible field $\boldsymbol{h}_{\text{rev}}$, which has to be stored as an internal variable during the numerical simulation.
 Note that by contrast with ferromagnetic hysteresis with a nonlinear saturation law [27, 29], in this context of superconducting hysteresis, the relationship Eq.
@@
 4.2 CS cell
-For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. We define
+For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. With
 $\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}},$ (28)
-and denote by $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
+and $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
 $\boldsymbol{g}=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (29)
@@
 $\approx\boldsymbol{h}_{\text{rev}}^{\text{trial}}+\frac{\tau_{\text{e}}+\tau_{\text{c}}}{\Delta t}(\boldsymbol{h}_{\text{rev}}^{\text{trial}}-\boldsymbol{h}_{\text{rev(p)}}),$ (32)
-with $\boldsymbol{h}_{\text{rev(p)}}$ the reversible field at the previous time step and $\Delta t$ the current time step.
+with $\boldsymbol{h}_{\text{rev(p)}}$ the reversible field at the previous time step and $\Delta t$ the current time step size.
 Solving this equation for $\boldsymbol{h}_{\text{rev}}^{\text{trial}}$, and then evaluating $\boldsymbol{h}_{\text{coupling}}^{\text{trial}}=\tau_{\text{c}}\dot{\boldsymbol{h}}_{\text{rev}}^{\text{trial}}$ gives the trial value (we drop the $\approx$ sign)
@@
 $\boldsymbol{h}_{\text{rev}}$ $=\boldsymbol{g}-\boldsymbol{h}_{\text{coupling}}-\boldsymbol{h}_{\text{eddy}}.$ (38)
-All instantaneous power quantities can be computed by Eqs. (12), (13), (17), (20), and (21).
+The instantaneous power quantities can be computed by Eqs. (12), (13), (18), (21), and (22).
 In the case of field-dependent irreversibility parameters $\kappa$ and $\chi$, a good convergence is again obtained in a few fixed points iterations, as for the S cell.
-The simulation of the CS cell requires to save the values of vectors $\boldsymbol{g}$ and $\boldsymbol{h}_{\text{rev}}$ at each time step.
+The implementation of the CS cell is straightforward.
+It only consists in two tests and a few update equations.
+The simulation of the CS cell requires to keep track of the values of vectors $\boldsymbol{g}$ and $\boldsymbol{h}_{\text{rev}}$ at the previous time step.
 4.3 Chain of cells
-In a hysteresis model with a chain of cells as proposed in Section 3.5, each cell $k$ is subject to the same magnetic field $\boldsymbol{h}$ and produces a distinct magnetic flux density $\boldsymbol{b}_{k}$.
-The total magnetic flux density $\boldsymbol{b}$ is computed as a weighted sum of the $\boldsymbol{b}_{k}$ as defined in Eq.
+In a ROHM model with a chain of cells as proposed in Section 3.4, each cell $k$ is subject to the same magnetic field $\boldsymbol{h}$ and produces a distinct magnetic flux density $\boldsymbol{b}_{k}$.
+The total magnetic flux density $\boldsymbol{b}$ is then computed as a weighted sum of the $\boldsymbol{b}_{k}$, as defined in Eq.
 (26).
@@
 Let us consider a numerical domain $\Omega$.
-It is decomposed into the superconducting strand (or any other superconducting system) to homogenize, denoted as $\Omega_{\text{m}}$, and the complementary domain, denoted as $\Omega_{\text{m}}^{\text{C}}$.
+It is decomposed into the superconducting strand (or any other superconducting conductor) to be homogenized, denoted as $\Omega_{\text{m}}$, and the complementary domain, denoted as $\Omega_{\text{m}}^{\text{C}}$.
 For simplicity, we assume that $\Omega_{\text{m}}^{\text{C}}$ is non-conducting, but the approach can be easily extended to conducting domains as well.
@@
 […]
-These equations can be solved numerically with the finite element method, on a mesh, i.e., a spatial discretization of the numerical domain $\Omega$.
-As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is more natural to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [41, 42].
+These equations can be solved numerically with the FE method, on a mesh, i.e., a spatial discretization of the numerical domain $\Omega$.
+As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is easier to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [39, 40].
 In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, this formulation can be reduced to a $\phi$-formulation [29].
-Starting from an initial solution, it consists of finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
+Starting from an initial solution, it consists in finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
 […]
-with $\Phi(\Omega)$ and $\Phi_{0}(\Omega)$ appropriate function spaces, and where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$.
+where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$.
+The space $\Phi(\Omega)$ contains functions in $H^{1}(\Omega)$ that satisfy appropriate essential boundary conditions.
+The space $\Phi_{0}(\Omega)$ is the same space but with homogeneous essential boundary conditions.
+These function spaces are discretized with lowest-order node functions [41].
 The presence of the hysteresis law in Eq.
 (40) makes the system history-dependent.
-The fields $\boldsymbol{h}_{\text{rev}}$ and $\boldsymbol{g}$ must therefore be saved at each steps.
-They are defined in $\Omega_{\text{m}}$ only and are chosen to be element-wise constant vectors.
+The fields $\boldsymbol{h}_{\text{rev}}$ and $\boldsymbol{g}$ must therefore be saved at each time step.
+They are defined in $\Omega_{\text{m}}$ only and are chosen to be element-wise constant vectors, with no continuity constraint across element interfaces.
 The hysteresis law also makes the system nonlinear.
 An iterative scheme is therefore necessary for a numerical simulation.
-In this case, we observed that the Newton-Raphson method leads to efficient simulations.
+We observed that the Newton-Raphson method leads to very efficient simulations.
 From an initial iterate $\boldsymbol{h}_{0}$, it consists in solving successively a linearized version of Eq.
 (40) until a given convergence criterion is met.
-At iteration $i\in\mathbb{N}_{0}$, the formulation reads, in terms of the unknown field $\boldsymbol{h}_{i}$,
+At iteration $i$, the formulation reads, in terms of the unknown field $\boldsymbol{h}_{i}$, and the previous (known) iterate $\boldsymbol{h}_{i-1}$,
 […]
@@
 5 Parameter Identification
-The hysteresis model presented in Section 3 provides a flexible tool able to capture a variety of different hysteretic responses thanks to the approach with several cells connected as a chain.
-The parameter values of each of these cells must be properly chosen in order to reproduce reference solutions or measurement data.
+The ROHM model presented in Section 3 provides a flexible tool able to capture a variety of different hysteretic responses thanks to the approach with several cells connected as a chain.
+The parameter values of each of these cells must be properly chosen in order to reproduce reference solutions.
 In this section, we propose a simple identification procedure for chains of S cells and explain a heuristic for chains of CS cells, which will be further illustrated in Section 7.
@@
 5.1 Chain of S cells
-In a chain of S cells, there are $2N$ parameters to be defined: the constant weights $\alpha_{k}$ and the irreversibility parameters $\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa}_{k}$.
-The number of cells $N$ must also be chosen as a trade-off between computational cost, implementation easiness and accuracy.
+In a chain of S cells, there are $2N$ parameters to be defined: the irreversibility parameters $\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\bar{\kappa}_{k}$ and the constant weights $\alpha_{k}$.
+The number of cells $N$ must be chosen as a trade-off between computational cost, implementation easiness and accuracy.
 As a reference solution, we consider the magnetization curve of a superconducting strand subject to a unidirectional field $\boldsymbol{h}_{\text{app}}(t)=h_{\text{app}}(t)\hat{\boldsymbol{e}}_{y}=h_{\text{max}}\sin(2\pi ft)\hat{\boldsymbol{e}}_{y}$, with a sufficiently large field $h_{\text{max}}$ for the system to be fully penetrated, and at a sufficiently low frequency $f$ for the coupling and eddy current effects to be negligible.
-We compute the associated average magnetic flux density in the strand $\boldsymbol{b}_{\text{in}}$ as
+Based on the reference solution, we compute the average magnetic flux density in the strand $\boldsymbol{b}_{\text{in}}$ as
 $\boldsymbol{b}_{\text{in}}$ $=\mu_{0}\boldsymbol{h}_{\text{in}}+\mu_{0}\boldsymbol{m}=\mu_{0}\left(\boldsymbol{h}_{\text{app}}+\frac{1}{2}\boldsymbol{m}\right),$ (42)
@@
 (1) and (7).
-An example of such a curve for the composite strand described in Section 2 is given in Fig.
-15, where $\mu_{0}h_{\text{max}}=2$ T.
-The curve is obtained with a standard 2D $h$-$\phi$-formulation that does not model coupling currents between the filaments, at $f=0.01$ Hz.
+An example of the $\boldsymbol{b}_{\text{in}}$ and magnetization curves for the composite strand described in Section 2 is given in Fig.
+14, with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz.
+The solution is obtained with a standard 2D FE $h$-$\phi$-formulation without any filament coupling.
 Figure 14: Reference magnetization curve for the identification of a chain of S cells.
 […]
-Step 1: Fitting the field-dependent scaling
+Step 1: Fitting the field-dependent scaling.
 The first step is to fit the irreversibility parameter scaling, $f_{\kappa}(\boldsymbol{b})$, introduced in Eq.
@@
 In Fig.
-15, this corresponds to the purple curves labelled as Step 1.
-For a chain of S cell, using Eqs. (8) and (23), the total magnetization is given by
+14, this corresponds to the purple curves labelled as Step 1.
+For a chain of S cells, using Eqs. (8) and (26), the total magnetization is given by
 $\boldsymbol{m}$ $=\boldsymbol{b}/\mu_{0}-\boldsymbol{h}=\sum_{k=1}^{N}\alpha_{k}\boldsymbol{h}_{\text{rev},k}-\boldsymbol{h}$
@@
 $|m|=\sum_{k=1}^{N}\alpha_{k}\kappa_{k}(\boldsymbol{b})=f_{\kappa}(\boldsymbol{b})\underbrace{\sum_{k=1}^{N}\alpha_{k}\bar{\kappa}_{k}}_{m_{\text{max}}},$ (44)
-with $m_{\text{max}}$ (in A/m) the maximum magnetization. As a result, the scaling function $f_{\kappa}(\boldsymbol{b})$ directly describes the shape of the purple curves in Fig. 15.
-A simple approach consists in choosing $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ as a scaling over the whole field range.
-This however does not lead to very accurate results at low fields for which $j_{\text{c}}$ strongly depends on $\boldsymbol{b}$.
-The reason is that the strand magnetization depends on the actual field distribution inside it, which is not uniform, and not only on the average vector $\boldsymbol{b}_{\text{in}}$.
-The larger the filaments in the strand, the larger the variation of $\boldsymbol{b}$ within each filament.
-A better approach is to directly identify the scaling $f_{\kappa}(\boldsymbol{b})$ on the reference curve.
-The purple curves can therefore be used to directly define $f_{\kappa}(\boldsymbol{b})$ in the available field range.
-Outside of the field range, $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ can be used as a first approximation.
+with $m_{\text{max}}$ (A/m) the maximum magnetization.
+Therefore, the scaling function $f_{\kappa}(\boldsymbol{b})$ directly describes the shape of the purple curves in Fig.
+14, scaled by $m_{\text{max}}$.
+As a result, the scaling $f_{\kappa}(\boldsymbol{b})$ can be defined as an interpolation function of the purple curves, scaled by $m_{\text{max}}$.
+Outside of the field range of the reference solution, $f_{\kappa}(\boldsymbol{b})$ can be extrapolated smoothly, e.g., using the field-dependent critical current density as a scaling.
+The fact that the model reproduces a maximum magnetization equal to $m_{\text{max}}$ will be ensured in step 2.
 In some circumstances, the maximum magnetization may not be observed at $\boldsymbol{b}_{\text{in}}=\boldsymbol{0}$ exactly.
 In such a case, the scaling $f_{\kappa}$ can be written in terms of a weighted sum such as $u\boldsymbol{b}+(1-u)\mu_{0}\boldsymbol{h}$, with $u\in[0,1]$ to be chosen.
-Step 2: Choosing the $\bar{\kappa}_{k}$ and fitting the weights $\alpha_{k}$
+Note that choosing $f_{\kappa}(\boldsymbol{b})=j_{\text{c}}(\boldsymbol{b})/j_{\text{c}}(\boldsymbol{0})$ as a scaling over the whole field range, instead of defining it based on the purple curve as proposed above, does not lead to very accurate results, especially at low fields.
+The reason is that the strand magnetization depends on the actual field distribution inside it, which is not uniform (but unknown for the homogenized ROHM model), and not only on the average vector $\boldsymbol{b}_{\text{in}}$.
+Step 2: Choosing the $\bar{\kappa}_{k}$ and fitting the weights $\alpha_{k}$.
 The second step consists in finding the remaining parameters, $\bar{\kappa}_{k}$ and $\alpha_{k}$, in order to reproduce transition branches between fully magnetized states, such as those labelled as Step 2 in Fig.
-15.
+14.
 The proposed procedure consists in choosing a priori the $N$ values $\bar{\kappa}_{k}$ and then fixing the weights $\alpha_{k}$ accordingly.
-For simplicity, we arrange the $\bar{\kappa}_{k}$ values in increasing order with respect to $k$ and start with $\bar{\kappa}_{1}=0$ A/m.
-We consider the virgin magnetization curve represented by the dotted curve in Fig.
-15.
+For simplicity, we arrange the $\bar{\kappa}_{k}$ values in increasing order with respect to $k$ and we start with $\bar{\kappa}_{1}=0$ A/m.
+We consider the virgin magnetization curve represented by the green curve starting at origin in Fig.
+14.
 Initially, $h_{\text{rev},k}=0$, $\forall k=1,\dots,N$.
-As the magnetic field $h_{\text{in}}$ progressively increases, it reaches the surface of the cell spheres one by one (see Fig.
-8).
+As the magnetic field $h_{\text{in}}$ increases, it reaches the surface of the cell spheres one by one (see Figs.
+7 and 15).
 This happens at successive threshold field values that depend on the $\bar{\kappa}_{k}$ constants and on the $f_{\kappa}(b_{\text{in}})$ scaling determined in Step 1.
-We denote these threshold fields as $h_{k}$.
+We denote these threshold fields as $h_{k}$, as illustrated in Fig.
+16.
+Figure 15: Illustration of the magnetic field and a chain of S cells with $N=4$ (the first cell is with $\kappa_{1}=0$ A/m), from virgin state to $\|\boldsymbol{h}\|=h_{4}$.
+Each sphere (circle) represents one cell.
+The center of the spheres are at positions $\boldsymbol{h}_{\text{rev},k}$, established by the history of $\boldsymbol{h}$, and their field-dependent radius is equal to $\kappa_{k}(\boldsymbol{b})$.
+The relative scale is respected compared to Fig.
+16.
+Variations of the magnetic field $\boldsymbol{h}$ that stay inside the smallest non-degenerated sphere produce no loss.
+[…]
 For $h<h_{k}$, we have $h_{\text{rev},j}=0$, $\forall j=k,\dots,N$, such that only the first $k-1$ cells contribute to the magnetic flux density $b$.
-The weights $\alpha_{k-1}$ can therefore be successively determined by forcing the solution of the hysteresis model to match the reference solution at each $h_{k}$, for $k=2,\dots,N$.
+The weights $\alpha_{k-1}$ can therefore be consecutively adjusted by forcing the solution of the ROHM model to match the reference solution at each $h_{k}$, for $k=2,\dots,N$.
+This adjustment can be automated with a simple iterative routine.
+The influence of the weights $\alpha_{k}$ on the ROHM model solution is illustrated in Fig.
+16.
 The last weight $\alpha_{N}$ is calculated so that all weights add up to one.
-The choice of the number of cells $N$ and the a priori distribution of values for the $\bar{\kappa}_{k}$ depend on the sought accuracy and application. This is discussed in Section 6.
+The choice of the number of cells $N$ and the a priori distribution of values for the $\bar{\kappa}_{k}$ depend on the desired accuracy and application. This is discussed in Section 6.
 5.2 Chain of CS cells
-For a chain of CS cells, there are $5N$ parameters to choose: $\alpha_{k}$, $\kappa_{k}$, $\tau_{\text{e},k}$, $\tau_{\text{c},k}$, and $\chi_{k}$, for $k=1,\dots,N$.
+For a chain of CS cells, there are $5N$ parameters to be defined: $\kappa_{k}$, $\alpha_{k}$, $\tau_{\text{e},k}$, $\tau_{\text{c},k}$, and $\chi_{k}$, for $k=1,\dots,N$.
 To identify all these parameters, reference solutions at different frequencies are necessary.
-A reference solution at a sufficiently low frequency for the eddy and coupling current effects to be negligible can be used exactly as described in the previous section to identify the $\alpha_{k}$ and $\kappa_{k}$ (with both the $f_{\kappa}(\boldsymbol{b})$ scaling and the $\bar{\kappa}_{k}$ values).
+We decompose the identification procedure in three steps.
+Step 1: Reproducing the rate-independent regime.
+A reference solution at a sufficiently low frequency, so that the eddy and coupling current effects are negligible, can be used exactly as described in the previous section to define the $\kappa_{k}$ (with both the $f_{\kappa}(\boldsymbol{b})$ scaling and the $\bar{\kappa}_{k}$ values) and $\alpha_{k}$.
 This leads to a number of $3N$ remaining parameters.
-The time constants $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ can be chosen so that the positions of the peaks in eddy and coupling losses respectively correspond to those of reference solutions, such as shown in Fig.
+Step 2: Fitting the time constants.
+The time constants $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ can be chosen so that the positions of the peaks in eddy and coupling losses correspond to those of reference solutions, such as shown in Fig.
 3.
 Choosing identical values for all the cells already leads to good results within limited amplitude ranges.
-To reproduce the observation that the maxima in coupling and eddy loss are observed at lower frequencies when the field amplitude increases, one can choose larger values of $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ for cells associated with larger irreversibility parameters $\kappa_{k}$.
+To reproduce the observation that the maxima in coupling and eddy loss are observed at lower frequencies when the field amplitude increases, one can choose larger values of $\tau_{\text{e},k}$ and $\tau_{\text{c},k}$ for cells associated with larger irreversibility parameters.
 This will be illustrated in Section 7.
-Once the time constants are chosen, the only remaining parameters are the irreversibility parameters $\chi_{k}(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi}_{k}$.
-The scaling $f_{\chi}(\boldsymbol{b})$ can be chosen as in the static case (Step 1), based on a reference solution at a sufficiently high frequency for the dynamic effects to be visible in the magnetization curve, e.g., at $f=100$ Hz for the 54-filament strand, as shown in Fig.
-7.
+Step 3: Reproducing the coupled filament magnetization.
+Once the time constants are fixed, the only remaining parameters are the $\chi_{k}(\boldsymbol{b})=f_{\chi}(\boldsymbol{b})\bar{\chi}_{k}$.
+The scaling $f_{\chi}(\boldsymbol{b})$ can be defined as in the static case (Step 1), based on a reference solution at a sufficiently high frequency for the dynamic effects to be visible in the magnetization curve, e.g., at $f=100$ Hz for the 54-filament strand, as shown in Fig.
+6.
 Then, the values $\bar{\chi}_{k}$ can be identified in order to best reproduce the transition branches, as in Step 2 of the static case, but now with values $\bar{\chi}_{k}$ as unknowns instead of the weights $\alpha_{k}$ which are already fixed.
 6 Results with a chain of S cells
-(a) Major loop $h_{\text{app}}^{(1)}(t)$.
-(b) Minor loops $h_{\text{app}}^{(2)}(t)$.
-Figure 16: Comparison of the hysteresis model results with a reference solution.
-Unidirectional transverse magnetic field.
-Chain of S cells with $N=6$ cells, parameters of Table 2.
-(a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$.
-(b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$.
-(Up) Magnetization loop.
-(Middle) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section.
-(Down) Internal magnetic field, the input of the hysteresis model.
-The first application consists in the 54-filament strand presented in Section 2, subject to a transverse field varying sufficiently slowly for the eddy and coupling current effects to be negligible, that is, with frequencies of the order of $f=0.01$ Hz.
-For the reference solution, we assume a non-conducting matrix and use a classical $h$-$\phi$-formulation [41], so that coupling effects are completely removed.
+The first application consists of the 54-filament strand presented in Section 2, subject to a transverse field varying sufficiently slowly for the eddy and coupling current effects to be negligible.
+We choose a frequency $f=0.01$ Hz.
+For the reference solution, we assume a non-conducting matrix and use a classical $h$-$\phi$-formulation [39], so that coupling effects are not present.
 Because of the finite value of the power index $n=30$, the strand response is not truly rate-independent (this would only be the case at the limit $n\to\infty$).
@@
 The reference solution is the major loop represented in Fig.
-15.
-We choose $N=6$ and apply the two-step procedure described in Section 5.1 with $N$ equally spaced values $\mu_{0}\bar{\kappa}_{k}$ from 00 T to $0.75$ T.
-The obtained parameters are given in Table 2.
-$k$ $\alpha_{k}$ $\mu_{0}\bar{\kappa}_{k}$
+14.
+We choose $N=6$ and apply the two-step procedure described in Section 5.1 with $N$ equally spaced values $\mu_{0}\bar{\kappa}_{k}$ from 00 mT to $750$ mT.
+The obtained parameters are given in Fig.
+17.
+$k$ $\mu_{0}\bar{\kappa}_{k}$ $\alpha_{k}$
 […]
-$1$ $48.24$ 00
-$2$ $18.07$ $150$
-$3$ $16.99$ $300$
-$4$ $9.31$ $450$
-$5$ $5.16$ $600$
-$6$ $2.22$ $750$
-Table 2: Parameters of a chain of S cells with $N=6$, identified on the reference magnetization loop shown in Fig. 15.
-It is interesting to notice that the first S cell, which is anhysteretic, almost contributes to half of the magnetic flux density.
-This is due to the large fraction of normal conductor in the strand, the matrix, which does not behave as a hysteretic material.
+$1$ 00 $48.24$
+$2$ $150$ $18.07$
+$3$ $300$ $16.99$
+$4$ $450$ $9.31$
+$5$ $600$ $5.16$
+$6$ $750$ $2.23$
+Figure 17: (Left) Parameters of a chain of S cells with $N=6$, identified on the reference magnetization loop shown in Fig. 14. (Right) Scaling function $f_{\kappa}(b)$.
+It is interesting to notice that the first S cell, which is anhysteretic ($\kappa_{1}=0$ A/m), almost contributes to half of the magnetic flux density.
+This is due to the large fraction of normal conductor in the strand, the copper matrix (copper fraction in the cross section is $56\%$), which does not behave as a hysteretic material.
+(a) Major loop $h_{\text{app}}^{(1)}(t)$.
+(b) Minor loops $h_{\text{app}}^{(2)}(t)$.
+Figure 18: Comparison of the ROHM model results with the reference solution.
+Unidirectional transverse magnetic field.
+Chain of S cells with $N=6$ cells, parameters of Fig.
+17.
+(a) Harmonic excitation $h_{\text{app}}^{(1)}(t)$.
+(b) Biharmonic excitation $h_{\text{app}}^{(2)}(t)$.
+(Up) Internal magnetic field, the input of the ROHM model.
+(Middle) Magnetization loop.
+The curves of the ROHM model are colored as a function of the instantaneous power, see the bottom plots for the legend.
+(Down) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section.
 6.2 Unidirectional excitation
-Using the parameters obtained above, we compare the predictions of the hysteresis model in terms of magnetization and loss for two unidirectional excitations along $\hat{\boldsymbol{e}}_{y}$.
+Using the parameters obtained above, we compare the predictions of the ROHM model in terms of magnetization and loss for two types of unidirectional excitations along $\hat{\boldsymbol{e}}_{y}$.
 The first one is a harmonic field:
 $\mu_{0}h_{\text{app}}^{(1)}(t)$ $=\mu_{0}h_{\text{max}}\sin(2\pi ft),$ (45)
-with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. The second one is a biharmonic field, so as to produce minor magnetization loops:
+with $\mu_{0}h_{\text{max}}=2$ T and $f=0.01$ Hz. The second one is a biharmonic field, resulting in a response with minor magnetization loops:
 $\mu_{0}h_{\text{app}}^{(2)}(t)$ $=\mu_{0}h_{\text{max}}\left(\sin(\pi ft)+0.25\sin(6\pi ft)\right).$ (46)
 Results are given in Fig.
-16.
-With $N=6$ cells, the hysteresis model is capable of reproducing faithfully both major and minor magnetization loops.
-The loss is accurately reproduced in both cases, with a relative error on the total loss below $1\%$.
-The influence of the number of cells on the relative error on total loss is shown in Fig.
-17.
-For the considered excitations, increasing the number of cells helps reducing the error for low $N$ values ($\leq 3$), but the error then stabilizes to a non-zero value.
-One source of error is related to the scaling function $f_{\kappa}(\boldsymbol{b})$ for irreversibility parameters, which is only approximate.
-Indeed, in reality, the magnetic flux density $\boldsymbol{b}$ is not uniform in the filament, but rather varies due to the screening currents and geometrical effects.
-Also, as mentioned above, the reference solution is not truly rate-independent, due to the finite $n$ value in the power law.
-Figure 17: Evolution of the relative error $(Q_{\text{hyst}}-Q_{\text{ref}})/Q_{\text{ref}}$ between total losses predicted by the hysteresis model ($Q_{\text{hyst}}$) and the reference model ($Q_{\text{ref}}$), for different sources, as a function of the number of cells $N$, with equally spaced values $\mu_{0}\bar{\kappa}_{k}$, from 00 T to $0.75$ T.
+18.
+With $N=6$ cells, the ROHM model reproduces very well both major and minor magnetization loops.
+The instantaneous loss is remarkably well reproduced in both cases, with a relative error on the total loss below $1\%$.
+The influence of the number of cells on the relative error on the total loss is shown in Fig.
+19.
+Increasing the number of cells helps reducing the error for low $N$ values ($\leq 3$).
+The error then stabilizes (for $N\geq 6$) to a small, non-zero value.
+Figure 19: Evolution of the relative error $(Q_{\text{hyst}}-Q_{\text{ref}})/Q_{\text{ref}}$ between total losses predicted by the ROHM model ($Q_{\text{hyst}}$) and the reference model ($Q_{\text{ref}}$), for different excitations, as a function of the number of cells $N$, with equally spaced values $\mu_{0}\bar{\kappa}_{k}$, from 00 T to $0.75$ T.
 In the case $N=1$, $\mu_{0}\bar{\kappa}_{1}=0.168$ T and $\alpha_{1}=1$.
 The major advantage of the method is its computational efficiency.
-Once the model parameters are identified, solving the hysteresis model is tremendously faster than solving a finite element model (typically a few milliseconds compared to a few tens of minutes).
-This efficiency is crucial in view of homogenizing full-scale superconducting systems.
-In practice, the number of cells and the distribution of $\bar{\kappa}_{k}$ values must be chosen depending on the actual excitations to be considered in the end application.
-A chain of S cells produces no loss for field variations smaller than the smallest non-zero irreversibility parameter $\kappa_{k}(\boldsymbol{b})$, as illustrated in Fig.
-18.
+Once the model parameters are identified, solving the ROHM model is tremendously faster than solving a detailed FE model.
+Typically a few milliseconds only is required for the ROHM model compared to a few tens of minutes for the 54-filament reference FE simulation.
+This efficiency is crucial in view of simulating full-scale superconducting systems.
+In practice, the number of cells and the distribution of $\bar{\kappa}_{k}$ values must be chosen depending on the actual excitations to be considered in practice.
+As already mentioned, a chain of S cells produces no loss for field variations smaller than the smallest non-zero irreversibility parameter $\kappa_{k}(\boldsymbol{b})$, as illustrated in Fig.
+15.
 To reproduce accurately power loss at low fields, or for low field ripples, the chain of S cells must therefore contain cells with sufficiently small irreversibility parameters.
-This will be further illustrated in the case of a chain of CS cells in Section 7.
-Figure 18: Illustration of a chain of S cells with $N=3$.
-The centers of the spheres are at positions $\boldsymbol{h}_{\text{rev},k}$, established by the history of $\boldsymbol{h}$.
-Variations of the magnetic field $\boldsymbol{h}$ that stay inside the smallest sphere do not produce any loss.
+The same comment holds for a chain of CS cells, and will be further illustrated in Section 7.
 6.3 Bidirectional excitation
-The parameters were identified based on the results of a unidirectional excitation, but all the hysteresis model equations are vectorial, and the model is therefore directly applicable to general excitations.
+The parameters in Section 6.1 are defined based on the results of a unidirectional excitation, but all the ROHM model equations are vectorial, and the model is therefore directly applicable to general excitations.
 We show in Fig.
@@
 The influence of the number of cells $N$ on the relative error is shown in Fig.
-17, a constant error of $\approx 3\%$ remains for large values of $N$.
-Figure 19: Comparison of hysteresis model results with a reference solution.
+19, a constant error of $\approx 3\%$ remains for large values of $N$.
+Figure 20: Comparison of ROHM model results with the reference solution.
 Rotating transverse magnetic field $\boldsymbol{h}_{\text{app}}^{(3)}(t)$.
-Chain of S cells with $N=6$ cells and parameters of Table 2.
-Solutions at selected time instants are indicated by the colored circles (reference solution) and crosses (hysteresis model).
-(Up) Internal magnetic field and magnetization of the filament (same legend as in the bottom sub-figure).
+Chain of S cells with $N=6$ cells and parameters of Fig.
+17.
+Solutions at selected time instants are indicated by the colored circles (reference solution) and crosses (ROHM model).
+(Up) Internal magnetic field and magnetization of the filament.
+The ROHM model curve is colored as a function of the instantaneous power loss.
+Same legend as in the bottom sub-figure.
 Values are in tesla (T).
 (Down) Dissipated power per unit length as a function of time.
-The hysteresis model reproduces well the magnetization and power loss.
+The ROHM model reproduces well the magnetization and power loss.
 The angle between the inner field $\boldsymbol{h}_{\text{in}}$ and the magnetization vector $\boldsymbol{m}$ is faithfully described, as shown by the solutions at selected time instants in Fig.
-19.
-The agreement is not perfect, but very satisfying provided that the model parameters were fully identified using results from a unidirectional situation, with absolutely no information about the strand response in more general cases.
-We let the questions of improving the agreement and tackling anisotropic systems for further works.
+20.
+The agreement is not perfect, but very satisfying provided that the model parameters are fully identified using results from a unidirectional situation, with absolutely no information about the strand response under bidirectional excitations.
 7 Results with a chain of CS cells
-As a generalization of the previous section, we consider the same 54-filament strand subject to a transverse field, but now with higher rates of field changes, so that both magnetization and loss exhibit frequency dependence.
-The reference solutions are obtained as described in Section 2 with the CATI method in order to account for coupling current effects.
-We start the analysis by choosing the hysteresis model parameters based on reference solutions. We then illustrate the results for wide ranges of field amplitudes and frequencies.
+As a generalization of the previous section, we consider the same 54-filament strand subject to a transverse field, but now with a conducting matrix and higher rates of field changes, so that both magnetization and loss exhibit frequency dependence.
+The reference solutions are obtained as described in Section 2 with the CATI method in order to account for coupling current effects due to the twist of the filaments.
+We start the analysis by choosing the ROHM model parameters based on reference solutions. We then illustrate the results for wide ranges of field amplitudes and frequencies.
 7.1 Parameter identification
 We choose the chain structure represented in Fig.
-14.
+21.
 Compared to the chain of S cells of the previous section, two cells now have a zero irreversibility parameter, $\kappa_{1}=0$ A/m and $\kappa_{2}=0$ A/m.
-They allow to reproduce losses at low fields, for which filament loss is small with respect to coupling and eddy loss.
 These two cells allow to reproduce the transitions between the three regimes illustrated in Fig.
 22 at low fields, associated with different magnetization slopes, as shown in Fig.
-21.
+23.
 The curves are obtained with the parameter values of Table 1, which are discussed below.
-Figure 20: Field distribution at low fields ($\mu_{0}h=10$ mT here) and different frequencies.
+Figure 21: Chain of CS cells to model a multifilamentary strand rate-dependent response.
+Figure 22: Field distribution at low fields ($\mu_{0}h_{\text{max}}=10$ mT) and different frequencies.
 The lines are parallel to the local magnetic field, the larger and darker they are, the higher the field amplitude.
-(a) Uncoupled filament and limited magnetization.
+(a) Uncoupled filaments and limited magnetization.
 (b) Increased magnetization due to coupled filaments.
-(c) Close to perfect diamagnetism due to eddy currents.
+(c) Almost perfect diamagnetism due to eddy currents.
 Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity).
-Dashed lines are reference solutions and solid lines are from the hysteresis model.
+Dashed lines are reference solutions and solid lines are from the ROHM model.
 At low frequencies, the dominant contribution to magnetization and loss comes from superconducting filament hysteresis.
-We can reuse the material parameters found in the previous section for the chain of S cells as was given in Table 2.
-In order to reproduce losses at fields lower than $0.15$ T, we also introduce $8$ new cells with lower irreversibility parameters, logarithmically spaced from $1$ mT to $80$ mT.
-The associated weights can be identified as proposed in Section 5.1, or based on analytical solutions at low fields.
+We can reuse the material parameters found in the previous section for the chain of S cells as is given in Fig.
+17.
+In order to reproduce losses at fields lower than $150$ mT, we also introduce $8$ new cells with lower irreversibility parameters, logarithmically spaced from $1$ mT to $80$ mT.
+The associated weights can be identified automatically as proposed in Section 5.1, or based on analytical solutions at low fields.
+These additional cells are optional and only needed to reproduce accurate results on wide field amplitude ranges.
 We then choose the time constants.
@@
 3 and discussed in Section 2.3.
-Finally, we fix the scaling $f_{\chi}(\boldsymbol{b})$ and the parameters $\bar{\chi}_{k}$. We found that a decent fit is obtained with
-$f_{\chi}(\boldsymbol{b})=\frac{1-\|\boldsymbol{b}\|/b_{0}}{1+\|\boldsymbol{b}\|/b_{1}},$ (45)
-with $b_{0}=15$ T and $b_{1}=4$ T. The parameters $\bar{\chi}_{k}$ are then tuned manually to obtain the values in Table 3.
-$k$ $\alpha_{k}$ $\mu_{0}\bar{\kappa}_{k}$ $\tau_{\text{e},k}$ $\tau_{\text{c},k}$ $\mu_{0}\bar{\chi}_{k}$
+Finally, we fix the scaling $f_{\chi}(\boldsymbol{b})$ and the parameters $\bar{\chi}_{k}$.
+A decent fit is obtained with the curve in Fig.
+24.
+The parameters $\bar{\chi}_{k}$ are then tuned manually to obtain the values in Table 1.
+Figure 24: Field-dependent scaling functions for the irreversibility parameters $\kappa_{k}(b)$ and $\chi_{k}(b)$ for the chain of CS cells.
+$k$ $\mu_{0}\bar{\kappa}_{k}$ $\alpha_{k}$ $\tau_{\text{e},k}$ $\tau_{\text{c},k}$ $\mu_{0}\bar{\chi}_{k}$
 […]
-$1$ $16.38$ $0.0$ $0.02$ $0.00$ $0.00$
-$2$ $22.62$ $0.0$ $0.02$ $0.08$ $1.13$
-$3$ $0.05$ $1.0$ $0.02$ $0.08$ $0.01$
-$4$ $0.10$ $1.9$ $0.02$ $0.08$ $0.02$
-$5$ $0.19$ $3.5$ $0.02$ $0.08$ $0.03$
-$6$ $0.35$ $6.5$ $0.02$ $0.08$ $0.07$
-$7$ $0.66$ $12.2$ $0.02$ $0.08$ $0.09$
-$8$ $1.24$ $22.9$ $0.02$ $0.08$ $0.14$
-$9$ $2.32$ $42.9$ $0.02$ $0.08$ $0.21$
-$10$ $4.33$ $80.2$ $0.02$ $0.08$ $0.40$
-$11$ $18.07$ $150.0$ $0.30$ $0.28$ $1.50$
-$12$ $16.99$ $300.0$ $0.30$ $0.28$ $2.10$
-$13$ $9.31$ $450.0$ $0.30$ $0.28$ $3.15$
-$14$ $5.16$ $600.0$ $0.30$ $0.28$ $4.20$
-$15$ $2.22$ $750.0$ $0.30$ $0.28$ $5.25$
+$1$ $0.0$ $16.38$ $0.02$ $0.00$ $0.00$
+$2$ $0.0$ $22.62$ $0.02$ $0.08$ $1.13$
+$3$ $1.0$ $0.05$ $0.02$ $0.08$ $0.01$
+$4$ $1.9$ $0.10$ $0.02$ $0.08$ $0.02$
+$5$ $3.5$ $0.19$ $0.02$ $0.08$ $0.03$
+$6$ $6.5$ $0.35$ $0.02$ $0.08$ $0.07$
+$7$ $12.2$ $0.66$ $0.02$ $0.08$ $0.09$
+$8$ $22.9$ $1.24$ $0.02$ $0.08$ $0.14$
+$9$ $42.9$ $2.32$ $0.02$ $0.08$ $0.21$
+$10$ $80.2$ $4.33$ $0.02$ $0.08$ $0.40$
+$11$ $150.0$ $18.07$ $0.30$ $0.28$ $1.50$
+$12$ $300.0$ $16.99$ $0.30$ $0.28$ $2.10$
+$13$ $450.0$ $9.31$ $0.30$ $0.28$ $3.15$
+$14$ $600.0$ $5.16$ $0.30$ $0.28$ $4.20$
+$15$ $750.0$ $2.22$ $0.30$ $0.28$ $5.25$
 Table 1: Parameters of a chain of CS cells with $N=15$.
 7.2 Results and numerical performance
-Using the model parameters given in Table 3, we compare the predictions of the hysteresis model with the reference solutions.
-The power loss per cycle as a function of frequency and selected magnetization loops are given in Figs.
-22 and 23.
-The power loss per cycle is also presented as a map in Fig.
-24, by interpolating the results of $400$ simulations, spanning over $25$ frequencies and $16$ field amplitudes, logarithmically spaced.
-Figure 22: Power loss per cycle for three different field amplitudes as a function of frequency.
-Solid curves are results from the hysteresis model with parameters of Table 3 and dashed curves are reference solutions.
+Using the model parameters given in Table 1, we compare the predictions of the ROHM model with the reference solutions.
+The loss per cycle as a function of frequency and magnetization loops are given in Figs.
+25 and 26.
+Figure 25: Energy loss per cycle for three different field amplitudes as a function of frequency.
+Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions.
 The legend is the same for the three subfigures.
 Figure 26: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies.
-Solid curves are results from the hysteresis model with parameters of Table 3 and dashed curves are reference solutions.
-Figure 24: Total loss per cycle and per unit length as a function of the frequency and amplitude of the applied field.
-Dashed curves are contour lines of equal loss.
-Solid curves delimit areas of distinct dominant loss contributions (from left to right: hysteresis in uncoupled filaments, coupling loss, hysteresis in coupled filaments, and eddy losses).
-(Left) Reference model, adapted from [31].
-(Right) Prediction of the hysteresis model with the parameters given in Table 3.
+Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions.
 The overall agreement on the total loss and magnetization is very good.
-The hysteresis model correctly reproduces the different regimes of distinct dominant loss contributions and describes the magnetization faithfully.
-The difference with the reference solution is higher at high frequency and high fields, as can be seen on the contour lines in the top right corners in Fig.
-24.
-This can be improved with a careful choice of model parameters, but as discussed in Section 3.3, the matching will never be perfect in the limit $f\to\infty$ as the CS cells do not reproduce a $1/\sqrt{f}$ decrease of the loss per cycle, but rather a $1/f$ decrease.
-For not too high frequencies however, the importance of this effect is limited.
-The major advantage of the hysteresis model compared to the detailed simulations is the very small associated computational work.
-Even with $N=15$ cells, the whole loss map on the right of Fig.
-24, which consists of $400$ hysteresis model simulations, can be performed in a few seconds, whereas the reference model requires $26$ days of computation time in total to generate the loss map on the left of Fig.
-24.
+The ROHM model correctly reproduces the different regimes of distinct dominant loss contributions and describes the magnetization faithfully.
+The difference with the reference solution is larger at high frequency and high fields, as can be seen in Fig.
+25.
+This can be improved with an adapted choice of model parameters, but as discussed in Section 3.3, the matching will never be perfect in the limit $f\to\infty$ as the CS cells do not reproduce a $1/\sqrt{f}$ decrease of the loss per cycle, but rather a $1/f$ decrease.
+For not too high frequencies however (skin depth not too small compared to strand diameter), the importance of this effect is limited.
+As in the rate-independent case, the major advantage of the ROHM model compared to the detailed simulations is the very small associated computational work.
+Even with $N=15$ cells, each ROHM model simulation is conducted in a few milliseconds, whereas the computational time for a reference solution with the CATI method is of the order of one hour (with 28 0002800028\,00028 000 degrees of freedom).
+The ROHM model enables the simulation of large-scale superconducting systems involving hundreds or thousands of strands, which would be completely unattainable with detailed strand models.
 […]
@@
 In superconducting magnets, many strands are combined together into cables, which themselves are wound around the magnet aperture.
-The hysteresis model can be used to homogenize the large number of strands into an equivalent homogeneous material.
-In such a case, which is let for further work, the reference detailed solution should be that of a close-packed set of strands, which can typically be modelled via appropriate periodic boundary conditions [14], rather than a single strand in air, as was done here for the sake of illustration.
-This may result in different model parameters, but the general approach is unchanged.
+The ROHM model can be used to homogenize the large number of strands into an equivalent homogeneous material.
+In such a case, the reference detailed solution should be that of a close-packed set of strands, which can typically be modelled via appropriate periodic boundary conditions [14], rather than a single strand in air, as is done here for the sake of illustration.
+This may result in different model parameters, but the general approach remains the same.
 8 Conclusion
-In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous loss in composite superconductors subject to transient external fields.
-This model is designed by adapting to superconducting systems the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27].
+In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous power loss in composite superconductors, subject to transient transverse external fields.
+This model is designed by adapting the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27, 28].
 We focused specifically on the case of a multifilamentary superconducting strand and proposed two models: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects.
 We proposed a parameter identification approach for both models and described their inclusion in a finite element framework.
-Finally, we demonstrated that these models help to strongly reduce the computational cost compared to conventional simulations, while offering a very good accuracy in a wide range of field amplitudes and rates of field change.
+Finally, we demonstrated that these models have the potential to reduce the computational cost compared to conventional simulations by several orders of magnitude.
+The model accuracy can be tailored for specific applications.
+A very good accuracy (typically with a few percent error on the power loss) in a wide range of field amplitudes and rates of field change can be obtained if required.
 This work constitutes a first step towards the homogenization of superconducting magnets, in which the large number of turns makes detailed simulations unrealistically expensive.
 Replacing the coil windings by homogenized material properties such as the proposed ROHM model allows to describe magnetization and loss efficiently.
-Further work is necessary to include non-zero transport current, which was not covered in this paper.
-Extension of the method to anisotropic systems such as high-temperature superconducting tapes or stacks of tapes in view of their homogenization can also be considered as further work.
+Further work is necessary to include non-zero transport current, which is not covered in this paper.
+Extension of the method to anisotropic systems such as high-temperature superconducting (HTS) tapes, stacks of HTS tapes, or HTS bulks in view of their homogenization can also be considered as further work.
 Appendix A Magnetization in 2D Problems
 In this section, we justify the introduction of a factor $2$ in the magnetization Eq. (1), in the case of an infinitely long problem solved in 2D with perpendicular currents.
-The magnetic dipole moment $\boldsymbol{M}$ of a current density distribution $\boldsymbol{j}$ in a given volume $V$ is evaluated as
+The magnetic dipole moment $\boldsymbol{M}$ (A m2) of a current density distribution $\boldsymbol{j}$ in a given volume $V$ is evaluated as
 $\boldsymbol{M}=\frac{1}{2}\int_{V}\boldsymbol{x}\times\boldsymbol{j}\,\text{d}V,$ (48)
@@
 Consider a rectangular current tube of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig.
-26.
+28.
 Accounting for the four sides of the rectangle, the contribution of this current loop to the total magnetic dipole moment is $wl\,\text{d}I$.
-Long and short sides of the rectangle contribute equally to the total magnetic dipole moment.
-Figure 26: Magnetic dipole moment of a current tube in a cylinder.
-When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D and per unit length, on a cross section of the conducting cylinder.
-In such a case, however, one must still account for the fact that all current loops close at the end of the cylinder, i.e., at infinity, for computing the magnetization per unit length.
-These closing currents are not part of the 2D solution, and are then missing from any magnetization calculation performed solely on the 2D cross section.
-Because the closing currents contribute to exactly half of the magnetic dipole moment, one can account for them by introducing a factor $2$, as is done in Eq.
+The two long sides of length $l$ contribute to half of it: $wl\,\text{d}I/2$.
+The two short sides of length $w$, closing the current loop, contribute to exactly the other half of it: $wl\,\text{d}I/2$.
+Figure 28: Current carrying rectangular tube in a cylinder.
+When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D, per unit length, on a cross section of the conducting cylinder.
+To compute the magnetic dipole moment in such a case, however, one must still account for the fact that all current loops must close at the end of the cylinder.
+These closing currents are not part of the 2D solution, and are then missing from any calculation performed solely on the 2D cross section.
+Because the closing currents conveniently contribute equally to the magnetic dipole moment compared to the other currents, one can account for them by introducing a factor $2$, as in Eq.
 (1).
 Appendix B Applied Field and Internal Field
-The power loss per cycle $Q_{\text{tot}}$ given by Eq.
-(4) can equivalently be computed in terms of $\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\boldsymbol{m}/2$.
+The energy loss per cycle $Q_{\text{tot}}$ given by Eq.
+(4) can be equivalently computed in terms of $\boldsymbol{h}_{\text{in}}=\boldsymbol{h}_{\text{app}}-\boldsymbol{m}/2$.
 Indeed, as also shown in [36],
@@
 $=a\oint\mu_{0}\boldsymbol{h}_{\text{in}}\cdot\text{d}\boldsymbol{m}.$ (49)
-That is, the areas enclosed by the loops represented in Figs. 5 and 7 are identical.
-Appendix C Hysteresis Model Jacobian
+In particular, this justifies that the areas enclosed by the loops represented in Figs. 4 and 6 are identical.
+Appendix C ROHM model Jacobian
 […]
 $\frac{\partial\boldsymbol{b}}{\partial\boldsymbol{h}}=\sum_{k=1}^{N}\alpha_{k}\mu_{0}\frac{\partial\boldsymbol{h}_{\text{rev},k}}{\partial\boldsymbol{h}}.$ (50)
-Its expression can be computed independently for each cell. Below, we give its expression for the simple case of a S cell and then generalize for a CS cell. We drop the index $k$ for conciseness.
+Its expression can be computed independently for each cell. Below, we give its expression for the simple case of a S cell and we then generalize for a CS cell. We drop the index $k$ for conciseness.
 C.1 S cell
@@
 $=\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}}),$ (51)
-with $\mathbf{I}$ the second-order identity tensor and $\mathbf{K}$ defined by the following dyadic product:
-$\mathbf{K}=(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}).$ (50)
+with $\mathbf{I}$ the second-order identity tensor and $\mathbf{K}$ the dyadic product $\mathbf{K}=(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}})$.
 C.2 CS cell
@@
 where $\boldsymbol{\mathcal{W}}_{\kappa}$ is the second-order tensor defined in Eq. (C.1), written here as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$.
-By contrast, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, one must use Eqs. (35), (33), and (34), which yields the following reversible field
+By contrast, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, one must use Eqs. (38), (36), and (37), which yield the following reversible field
 $\boldsymbol{h}_{\text{rev}}=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\left(\boldsymbol{g}-\chi\frac{\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|}+\frac{\tau_{\text{e}}}{\Delta t}\boldsymbol{h}_{\text{rev(p)}}\right)$ (54)
@@
 $+\frac{\chi}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|^{3}}\mathbf{L},$ (56)
-with
-$\mathbf{L}=(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}).$ (56)
+with $\mathbf{L}=(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})\otimes(\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}})$.

### v3

@@
 In this contribution, we focus on the case of a strand with twisted superconducting filaments subject to a time-varying transverse magnetic field.
-We propose two levels of ROHM models: (i) a rate-independent model that reproduces hysteresis in the filaments, and (ii) a rate-dependent model that generalizes the first level by also reproducing dynamic effects due to coupling and eddy currents.
-We describe the implementation and inclusion of these models in a finite element framework, discuss how to deduce the model parameters, and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency over a wide range of excitation frequencies and amplitudes.
-  • October 25, 2024
+We propose two variations of the ROHM model: (i) a rate-independent model that reproduces hysteresis in the filaments, and (ii) a rate-dependent model that generalizes the first level by also reproducing dynamic effects due to coupling and eddy currents.
+We then describe the implementation and inclusion of the ROHM model in a finite element framework, discuss how to deduce the model parameters, and finally demonstrate the capabilities of the approach in terms of accuracy and efficiency over a wide range of excitation frequencies and amplitudes.
+  • February 5, 2025
 Keywords: Reduced Order Method, Hysteresis Model, AC Loss, Magnetization.
 […]
-*                                    This work has been submitted to a journal for possible publication.
-Copyright may be transferred without notice, after which this version may no longer be accessible.
-[…]
-[…]
-*                                    This work has been submitted to a journal for possible publication.
-Copyright may be transferred without notice, after which this version may no longer be accessible.
 1 Introduction
-Large-scale superconducting magnets are multi-scale structures made of a high number of wound superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes.
-The magnetic response of these systems is hysteretic and rate-dependent.
-Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting elements [1, 2, 3, 4], whereas rate-dependency is caused by eddy and coupling current dynamics [5, 6, 7].
+Large-scale superconducting magnets are multi-scale structures made of multi-turn windings of superconducting cables, which themselves consist of multifilamentary strands or complex arrangements of tapes.
+The magnetic response of these systems is hysteretic and rate-dependent because of small-scale current density dynamics.
+Hysteresis is created by the irreversible motion of flux vortices among pinning centers in superconducting elements [1, 2, 3, 4].
+Rate-dependency is caused by eddy current and coupling current dynamics [5, 6, 7].
 Hysteresis, eddy current, and coupling current effects induce loss in transient regimes [5] as well as field distortions [8].
-Their accurate description and computation is therefore crucial, e.g., for computing the load on the cryogenic system, the temperature and stability margins, field errors, or for the design of quench protection devices.
+Their accurate description and computation is therefore crucial, e.g., for computing the load on the cryogenic system, temperature and stability margins, field errors, or for the design of quench protection devices.
 Most approaches to describe the superconducting hysteresis are based on the definition of a relationship between electric field and current density, either non-smooth, as in the Bean model [1], or smooth, as in the power law model [9].
 The use of the power law model leads to very accurate results and versatile numerical models with, e.g., the finite element (FE) method [10], which can also be combined with eddy current and coupling current models.
-Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes the numerical models computationally demanding to solve, already for single strands or cables [11].
+Besides these advantages, the power law introduces a strong nonlinearity in the equations, which makes numerical models of superconducting systems computationally demanding to solve, already for single strands or cables [11, 12, 13, 14].
 As a consequence, modelling large-scale superconducting systems in all their details down to the superconductor level is completely unrealistic in practice.
-A faster and more efficient method is necessary.
-Homogenization methods which reproduce the macroscopic effects of coupling dynamics without modelling them explicitly are good candidates.
-The general idea of these methods is to describe the systems in terms of the average fields [12], in order to reproduce the magnetization and loss accurately, without having to solve for the detailed current density distribution, which has the potential to strongly reduce the computational cost of the simulations.
-Such techniques have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures with bundles of wires of any shape [13, 14, 15], described in the frequency or time domain [16, 17].
-They rely on the design of equivalent conductivity and permeability parameters, that must be fitted on reference solutions.
-Following a similar approach for superconducting systems requires the use of dedicated history-dependent parameters in order to reproduce their hysteretic response.
+It is often necessary to group conductors together in order to lower simulation time [15, 16], and multi-scale procedures [17] or domain decomposition methods [13, 18] combined with parallel computing are being developed to speed up simulations.
+Homogenization methods describing systems in terms of the average fields [19] without modelling the small-scale field distributions explicitly are also good candidates for obtaining even more efficient large-scale superconducting magnet models.
+The general idea of these methods consists in replacing complex conducting structures by fictitious, uniform, homogenized materials with special properties.
+These properties are designed in order to reproduce directly the macroscopic effects of the small-scale current density distribution, e.g., magnetization and loss, without modelling that current density distribution explicitly.
+Once the homogenized material properties are defined by fitting their predictions with those of reference models, they can be used to reproduce accurately the response of large-scale systems with a strongly reduced computational cost compared to conventional detailed simulations.
+Such homogenization methods have been extensively studied in non-superconducting electromagnetic systems, such as periodical structures with bundles of wires of any shape [20, 21, 22], described in the frequency or time domain [23, 24].
+They rely on the definition of equivalent permeability and conductivity values (or functions), to account for proximity effect and skin effect, respectively.
+Following a similar approach for superconducting systems requires to define history-dependent material properties in order to reproduce a hysteretic response. This is the objective of this paper.
 A few hysteresis models have been proposed for modelling superconductors.
@@
 In 1997, Bergqvist proposed a thermodynamically consistent approach to model hysteresis [33], which led to the development of the energy-based model [34, 35, 36].
-The energy-based model offers several advantages.
+The energy-based model offers several advantages compared to other models.
 First, it gives a direct access to instantaneous dissipated and stored power by conveniently separating the magnetic field into several contributions.
 Second, it offers a high flexibility and modularity thanks to an approach involving several sub-elements that are combined to form the final model.
-Third, it is consistently defined as a vector, and not scalar model from the beginning [27], and hence applies to two-dimensional (2D) or three-dimensional (3D) settings.
+Third, it is consistently defined as a vector model from the beginning [34], and hence applies to two-dimensional (2D) or three-dimensional (3D) settings.
 Fourth, it can be directly generalized to include rate-dependent effects [35].
 Finally, the equations of this model consist in simple explicit tests, which leads to straightforward implementations and very efficient simulations.
-In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis by focussing on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis.
-This situation is relevant for most applications and constitutes a first step towards the homogenization of a complete magnet winding.
-This is also a challenging problem since, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects, eddy currents and interfilamentary coupling currents in the strand matrix, as well as filament hysteresis change due to coupling and eddy currents.
-For this paper, we assume that the strand carries no transport current.
-The inclusion of transport current will be the focus of further work.
+In this paper, we adapt and extend the state-of-the-art energy-based model to composite superconductor hysteresis, paying attention to preserving all the above-mentioned advantages.
+In this first contribution, we focus on the case of a superconducting multifilamentary strand subject to a transverse magnetic field, i.e., a field perpendicular to the strand axis, which is relevant for most magnet applications.
+The multifilamentary strand is also a challenging problem.
+Indeed, in addition to the filament hysteresis effect, it also involves two types of rate-dependent effects, interfilament coupling currents and eddy currents in the strand matrix, as well as filament hysteresis change due to coupling and eddy currents.
+In this paper, we assume that the strand carries no transport current.
+Accounting for transport current and the computing the associated voltage is necessary for the homogenization of a complete magnet winding, but will be the focus of a further work.
 We also consider a constant temperature and do not model thermal effects.
-We propose the Reduced Order Hysteretic Magnetization (ROHM) model, and present two variations of it: (i) a rate-independent model which describes the superconducting hysteresis, and (ii) a rate-dependent model, which generalizes the former by also describing eddy current and coupling current effects, in addition to the superconducting hysteresis.
-The ROHM model is general and can be adapted to a variety of superconductors or composite superconductors, such as strands, tapes, stacks of tapes, cables, or bulks.
-It can also be easily implemented in any FE framework that allows for user-defined material properties.
+[…]
+Although we apply it on multifilamentary strands only, the ROHM model can be adapted to a variety of superconductors or composite superconductors, such as strands, tapes, stacks of tapes, cables, or bulks.
 This paper is structured as follows.
 In Section 2, we describe the response of a multifilamentary strand in terms of magnetization and loss for transverse fields with a range of frequencies and amplitudes, in order to motivate the design of the ROHM model.
-In Section 3, we present the equations defining the rate-independent and rate-dependent ROHM models.
+In Section 3, we present the equations defining the rate-independent and rate-dependent variations of the ROHM model.
 In Section 4, we discuss the implementation of these equations and we describe how to include them in a 2D FE framework with a $\phi$-formulation.
 We then propose in Section 5 a parameter identification approach for fixing the ROHM model parameters based on reference solutions.
-Finally, in Sections 6 and 7, we apply the ROHM model in rate-independent and rate-dependent situations, respectively, and demonstrate its potential in the context of homogenization techniques.
+Finally, in Sections 6 and 7, we apply the ROHM model in rate-independent and rate-dependent situations, respectively, and demonstrate its potential as a homogenized model.
 2 Multifilamentary Strand Dynamics
-In this section, we describe the response of a composite superconducting strand subject to a transverse magnetic field in terms of power loss and magnetization.
+In this section, we describe the response of a composite superconducting strand subject to a transverse magnetic field in terms of the induced power loss and magnetization.
 To this end, we model the detailed current density distribution inside the strand with a FE method, the CATI method [37] implemented in FiQuS [38, 39, 40], and we analyze the different dynamics resulting from the strand composite structure.
@@
 2.3.1 Rate-independent regime
-At low frequencies (here for $f\lesssim 0.01$ Hz), the rate of change of the transverse magnetic field generates negligible coupling and eddy currents in the matrix.
-In this regime, the superconducting filaments behave almost uncoupled and most of the loss is due to hysteresis in them.
+At low frequencies (here for $f\lesssim 0.01$ Hz), the rate of change of the transverse magnetic field generates negligible coupling and eddy currents in the matrix [5, 17].
+In this regime, the superconducting filaments behave almost as uncoupled and most of the loss is due to hysteresis in them.
 A typical magnetization loop is shown in Fig.
@@
 27).
-However, to simulate transients in a full-scale magnet containing thousands of strands, one cannot afford models involving such a detailed description at the strand level.
+However, to simulate transients in a full-scale magnet cross section containing thousands of strands, one cannot afford models involving such a detailed description at the strand level.
 Instead, one can use an alternative model that produces accurate macroscopic response in terms of power loss and magnetization, which are both history-dependent and rate-dependent, but that does not rely on a detailed calculation of the current density.
 We present such a model in this section: the Reduced Order Hysteretic Magnetization (ROHM) model.
-This model is adapted from the energy-based model for ferromagnetic materials [27, 28].
-We start in Section 3.1 by clarifying how the ROHM model can be used to model the strand.
-We then define the building blocks, referred to as cells, that will be combined for the construction of the ROHM model.
-Each cell defines a local hysteretic relationship between the magnetic field $\boldsymbol{h}$ and the magnetic flux density $\boldsymbol{b}$.
+This model is inspired from the energy-based model for ferromagnetic materials [34, 35].
+We start in Section 3.1 by clarifying how the ROHM model can be used to model a superconducting strand.
+We then define the building blocks, referred to as cells, that will be used for the construction of the ROHM model.
+Each cell defines a local and elementary hysteretic relationship between the magnetic field $\boldsymbol{h}$ as an input, and the magnetic flux density $\boldsymbol{b}$ as an output.
 We present in Section 3.2 a rate-independent cell, that we call the superconductor cell (S cell).
-In Section 3.3, we add contributions to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell).
-Finally, we show in Section 3.4 how to combine several cells into a chain of cells, defining a complete ROHM model.
+In Section 3.3, we add components to account for eddy current and coupling current effects, resulting in a rate-dependent cell, that we call the composite superconductor cell (CS cell).
+Inside each cell, the magnetic field vector is expressed as the sum of distinct components: 2 components for the S cell, and 4 components for the CS cell.
+These components describe the interplay between different magnetization mechanisms and allow for a clear separation between stored and dissipated energies.
+A single cell is not sufficient for an accurate description of the strand response. We show in Section 3.4 how several cells can be combined into a chain of cells, defining the complete ROHM model.
 3.1 Homogenized model and internal field
@@
 3.3 Composite superconductor cell - CS cell
-For regimes in which the superconducting strand response is strongly rate-independent, the S cell presented in the previous section must be generalized.
+For regimes in which the superconducting strand response is strongly rate-dependent, the S cell presented in the previous section must be generalized.
 Starting from the S cell, we propose to account for the magnetization due to eddy currents and coupling currents by adding two new contributions in the magnetic field decomposition, $\boldsymbol{h}_{\text{eddy}}$ and $\boldsymbol{h}_{\text{coupling}}$, such that Eq.
@@
 For a magnetic field $\boldsymbol{h}$ at the new time step, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated as follows:
-$\boldsymbol{h}_{\text{rev}}$ =$&\boldsymbol{h}_{\text{rev(p)}},&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},&&\text{otherwise,}\end{aligned}\right.$
-$=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}}).$ (27)
+$\boldsymbol{h}_{\text{rev}}$ =$&\boldsymbol{h}_{\text{rev(p)}},&&\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},&&\text{otherwise,}\end{aligned}\right.$ (27)
 The associated magnetic flux density $\boldsymbol{b}$ is directly given by Eq.
@@
 In such a case, an approximate solution can be computed easily with a few fixed point iterations, that is, by solving successively Eq.
-(4.1) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
+(27) with updated values of $\kappa(\boldsymbol{b})=\kappa(\mu_{0}\boldsymbol{h}_{\text{rev}})$ until $\boldsymbol{h}_{\text{rev}}$ does no longer change significantly.
 The implementation of the S cell is very simple, as it entirely consists in the test of Eq.
-(4.1).
+(27).
 It only requires the knowledge of the previous reversible field $\boldsymbol{h}_{\text{rev}}$, which has to be stored as an internal variable during the numerical simulation.
@@
 The update rule Eq.
-(4.1) is therefore exact [29], up to the time discretization error.
+(27) is therefore exact [36], up to the time discretization error.
 4.2 CS cell
-For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements. With
-$\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}},$ (28)
-and $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step. We start by updating $\boldsymbol{g}$, as a function of the new $\boldsymbol{h}$ as follows:
-$\boldsymbol{g}=\boldsymbol{\mathcal{U}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (29)
-with the update rule defined by Eq.
-(4.1) but as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$.
-This operation contains the first test, after which we can already compute the irreversible field $\boldsymbol{h}_{\text{irr}}=\boldsymbol{h}-\boldsymbol{g}$.
+For the CS cell described in Section 3.3, the approach is similar. However, two tests are now necessary to account for the two dry friction elements.
+We start by updating $\boldsymbol{g}=\boldsymbol{h}_{\text{rev}}+\boldsymbol{h}_{\text{eddy}}+\boldsymbol{h}_{\text{coupling}}$ (see Fig.
+10).
+Denoting $\boldsymbol{g}_{\text{(p)}}$ its value at the previous time step, we define the update rule for $\boldsymbol{g}$ as follows:
+$\boldsymbol{g}$ =$&\boldsymbol{g}_{\text{(p)}},\text{ if }\|\boldsymbol{h}-\boldsymbol{g}_{\text{(p)}}\|\leq\kappa\text{ and }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\boldsymbol{h}-\kappa\frac{\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|},\text{ otherwise,}\end{aligned}\right.$ (28)
+where the second necessary condition in the first case ensures that $\|\boldsymbol{h}_{\text{irr}}\|<\kappa$ is possible only if $\dot{\boldsymbol{b}}=\boldsymbol{0}$.
+From $\boldsymbol{g}$, we can compute $\boldsymbol{h}_{\text{irr}}=\boldsymbol{h}-\boldsymbol{g}$.
 We proceed with the second test, encoded in Eq.
@@
 As the hysteresis law is driven by the magnetic field $\boldsymbol{h}$, it is easier to consider an $h$-conform formulation, such as the $h$-$\phi$-formulation [46, 47].
-In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, this formulation can be reduced to a $\phi$-formulation [29].
-Starting from an initial solution, it consists in finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
+In this particular case with no net currents in $\Omega_{\text{m}}$ and a non-conducting $\Omega_{\text{m}}^{\text{C}}$, the $h$-$\phi$-formulation can be reduced to a $\phi$-formulation [36].
+The inclusion of the ROHM model in a $b$-conform formulation is less straightforward and is discussed in [48].
+Starting from an initial solution, the $\phi$-formulation consists in finding a magnetic field $\boldsymbol{h}=\textrm{{grad}}\,\phi$ with $\phi\in\Phi(\Omega)$ such that, at subsequent time instants, $\forall\boldsymbol{h}^{\prime}=\textrm{{grad}}\,\phi^{\prime}$ with $\phi^{\prime}\in\Phi_{0}(\Omega)$,
 […]
 where the notation $\left(\boldsymbol{v}\,,\boldsymbol{w}\right)_{\Omega}$ denotes the integral over $\Omega$ of the dot product of any two vector fields $\boldsymbol{v}$ and $\boldsymbol{w}$.
-The space $\Phi(\Omega)$ contains functions in $H^{1}(\Omega)$ that satisfy appropriate essential boundary conditions.
-The space $\Phi_{0}(\Omega)$ is the same space but with homogeneous essential boundary conditions.
+The function space $\Phi(\Omega)$ contains scalar fields in $H^{1}(\Omega)$ that satisfy appropriate essential boundary conditions.
+The function space $\Phi_{0}(\Omega)$ is the same function space but with homogeneous essential boundary conditions.
 These function spaces are discretized with lowest-order node functions [49].
@@
 (b) Minor loops $h_{\text{app}}^{(2)}(t)$.
-Figure 18: Comparison of the ROHM model results with the reference solution.
+Figure 18: Comparison of the ROHM model results with reference solutions.
 Unidirectional transverse magnetic field.
@@
 (Middle) Magnetization loop.
-The curves of the ROHM model are colored as a function of the instantaneous power, see the bottom plots for the legend.
+The curves of the ROHM model are colored as a function of the instantaneous power, according to the values in the bottom plots.
 (Down) Dissipated power per unit length as a function of time, equal to $a\,p_{\text{irr}}(t)$, with $a=\pi d^{2}/4$ the surface area of the strand cross section.
@@
 Figure 23: Strand magnetization for $\mu_{0}h_{\text{max}}=10$ mT at different frequencies, from virgin state at time $t=0$ to time $t=0.25/f$ (for better clarity).
-Dashed lines are reference solutions and solid lines are from the ROHM model.
+Dashed curves are reference solutions.
+Solid curves are from the ROHM model.
 At low frequencies, the dominant contribution to magnetization and loss comes from superconducting filament hysteresis.
@@
 Figure 25: Energy loss per cycle for three different field amplitudes as a function of frequency.
-Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions.
+Solid curves are results from the ROHM model with parameters of Table 1.
+Dashed curves are reference solutions.
 The legend is the same for the three subfigures.
 Figure 26: Strand magnetization for $\mu_{0}h_{\text{max}}=1$ T at different frequencies.
-Solid curves are results from the ROHM model with parameters of Table 1 and dashed curves are results from the reference solutions.
+Solid curves are results from the ROHM model with parameters of Table 1.
+Dashed curves are reference solutions.
 The overall agreement on the total loss and magnetization is very good.
@@
 This may result in different model parameters, but the general approach remains the same.
-8 Conclusion
-In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous power loss in composite superconductors, subject to transient transverse external fields.
-This model is designed by adapting the state-of-the-art energy-based model developed in the context of ferromagnetic hysteresis [27, 28].
-We focused specifically on the case of a multifilamentary superconducting strand and proposed two models: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects.
-We proposed a parameter identification approach for both models and described their inclusion in a finite element framework.
-Finally, we demonstrated that these models have the potential to reduce the computational cost compared to conventional simulations by several orders of magnitude.
-The model accuracy can be tailored for specific applications.
-A very good accuracy (typically with a few percent error on the power loss) in a wide range of field amplitudes and rates of field change can be obtained if required.
-This work constitutes a first step towards the homogenization of superconducting magnets, in which the large number of turns makes detailed simulations unrealistically expensive.
-Replacing the coil windings by homogenized material properties such as the proposed ROHM model allows to describe magnetization and loss efficiently.
-Further work is necessary to include non-zero transport current, which is not covered in this paper.
-Extension of the method to anisotropic systems such as high-temperature superconducting (HTS) tapes, stacks of HTS tapes, or HTS bulks in view of their homogenization can also be considered as further work.
+8 Conclusion and outlook
+In this work, we introduced the Reduced Order Hysteretic Magnetization (ROHM) model to represent the magnetization and instantaneous power loss induced in composite superconductors by transient transverse magnetic fields.
+This model is designed by adapting the state-of-the-art energy-based model for ferromagnetic hysteresis [34, 35].
+We focused specifically on the case of a multifilamentary superconducting strand and proposed two variations of the ROHM model: a rate-independent model, describing hysteresis in superconducting filaments, and a rate-dependent model, accounting for filament coupling, coupling current, and eddy current effects.
+We proposed a parameter identification approach for both variations of the model, based on reference solutions obtained with detailed strand models, and we described their inclusion in a finite element framework.
+Finally, we demonstrated that the ROHM model has the potential to reduce the computational cost compared to conventional simulations by several orders of magnitude.
+The model accuracy can be tailored for specific applications: a very good accuracy (typically with a few percent error on the power loss) in a wide range of field amplitudes and rates of change can be obtained if required.
+It also allows to represent deformed strands or strands with defects, provided that reference solutions are used in order to fix adapted ROHM model parameters.
+This work constitutes a first step towards the homogenization of superconducting magnet windings, in view of enabling the simulation of magneto-thermal transients in superconducting magnet cross sections [50], accounting for all loss contributions.
+Replacing the cables with homogenized materials allows to describe the magnet response accurately and efficiently without computing the detailed current density distribution.
+The ROHM model accounts for magnetization and loss induced by an external transverse field.
+One next step consists in defining a second reduced order model to describe the inductance and power loss induced by transport current.
+The response of a composite strand to transport current is also hysteretic and rate-dependent, therefore, this second reduced order model can be built using similar concepts as those of the ROHM model.
+Additionally, the ROHM model and the reduced order model for transport current must be coupled.
+Indeed, the presence of a transport current affects the magnetization, and vice-versa, the presence of a background magnetic field affects the transport current response.
+Temperature dependence must also be included in the model parameters for magneto-thermal simulations.
+For low-temperature superconducting magnets, interstrand coupling current effects in Rutherford cables must also be accounted for.
+They can be studied using detailed cable models such as in [51] combined with homogenized strand models.
+The homogenization of these effects into reduced order models will be considered as another next step.
+The extension of the approach to high-temperature superconducting (HTS) tapes, stacks of HTS tapes, or HTS bulks in view of their homogenization requires to account for the strong anisotropy of these systems.
+In principle, as the ROHM model equations are all vectorial, this can be done via the definition of direction-dependent irreversibility parameters and time constants.
+Further study is required in that direction.
 Appendix A Magnetization in 2D Problems
-In this section, we justify the introduction of a factor $2$ in the magnetization Eq. (1), in the case of an infinitely long problem solved in 2D with perpendicular currents.
+In this section, we justify the introduction of a factor $2$ in the magnetization Eq. (1), in the case of an infinitely long problem solved in 2D with axial currents.
 The magnetic dipole moment $\boldsymbol{M}$ (A m2) of a current density distribution $\boldsymbol{j}$ in a given volume $V$ is evaluated as
@@
 with $\boldsymbol{x}$ the position vector.
-Consider a rectangular current tube of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig.
+Consider a rectangular current loop of cross section $\text{d}S$, width $w$, and length $l$, carrying a current $\text{d}I=\boldsymbol{j}\cdot\text{d}\boldsymbol{S}$, as illustrated in Fig.
 28.
@@
 The two short sides of length $w$, closing the current loop, contribute to exactly the other half of it: $wl\,\text{d}I/2$.
-Figure 28: Current carrying rectangular tube in a cylinder.
+Figure 28: Rectangular current loop in a cylinder.
 When the geometry is sufficiently long for end effects to be neglected, the problem is usually solved in 2D, per unit length, on a cross section of the conducting cylinder.
 To compute the magnetic dipole moment in such a case, however, one must still account for the fact that all current loops must close at the end of the cylinder.
-These closing currents are not part of the 2D solution, and are then missing from any calculation performed solely on the 2D cross section.
-Because the closing currents conveniently contribute equally to the magnetic dipole moment compared to the other currents, one can account for them by introducing a factor $2$, as in Eq.
+These closing currents are not part of the 2D solution, and are therefore missing from any calculation performed solely on the 2D cross section.
+Because the closing currents conveniently contribute equally to the magnetic dipole moment compared to the other currents, one can account for them by simply introducing a factor $2$, as in Eq.
 (1).
@@
 C.1 S cell
-For a S cell, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated with Eq. (4.1). The derivative of this update rule is given by
+For a S cell, the reversible field $\boldsymbol{h}_{\text{rev}}$ is updated with Eq. (27). The derivative of this update rule is given by
 $\frac{\partial\boldsymbol{h}_{\text{rev}}}{\partial\boldsymbol{h}}$ =$&\mathbf{0},\quad\text{if }\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|\leq\kappa,\\ &\left(1-\frac{\kappa}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|}\right)\mathbf{I}+\frac{\kappa}{\|\boldsymbol{h}-\boldsymbol{h}_{\text{rev(p)}}\|^{3}}\mathbf{K},\\ &\qquad\text{otherwise,}\end{aligned}\right.$
@@
 $\frac{\partial\boldsymbol{h}_{\text{rev}}}{\partial\boldsymbol{h}}$ $=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\frac{\partial\boldsymbol{g}}{\partial\boldsymbol{h}}$
-$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (53)
-where $\boldsymbol{\mathcal{W}}_{\kappa}$ is the second-order tensor defined in Eq. (C.1), written here as a function of $\boldsymbol{g}_{\text{(p)}}$ instead of $\boldsymbol{h}_{\text{rev(p)}}$.
+$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}+\tau_{\text{c}}}\tilde{\boldsymbol{\mathcal{W}}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}},\boldsymbol{g}_{\text{(p)}}),$ (52)
+where $\tilde{\boldsymbol{\mathcal{W}}}_{\kappa}$ is a second-order tensor similar to that defined in Eq.
+(C.1), but with $\|\boldsymbol{h}-\boldsymbol{g}_{\text{(p)}}\|\leq\kappa$ as an additional necessary condition for the first case, as in Eq.
+(28).
 By contrast, if $\|\boldsymbol{h}_{\text{coupling}}^{\text{trial}}\|>\chi$, one must use Eqs. (37), (35), and (36), which yield the following reversible field
@@
 $\frac{\partial\boldsymbol{h}_{\text{rev}}}{\partial\boldsymbol{h}}$ $=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\left(\frac{\partial\boldsymbol{g}}{\partial\boldsymbol{h}}-\frac{\partial}{\partial\boldsymbol{g}}\left(\chi\frac{\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}}{\|\boldsymbol{g}-\boldsymbol{h}_{\text{rev(p)}}\|}\right)\cdot\frac{\partial\boldsymbol{g}}{\partial\boldsymbol{h}}\right)$
-$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\boldsymbol{\mathcal{W}}_{\chi}(\boldsymbol{g},\boldsymbol{h}_{\text{rev(p)}})\cdot\boldsymbol{\mathcal{W}}_{\kappa}(\boldsymbol{h},\boldsymbol{g}_{\text{(p)}}),$ (55)
+$=\frac{\Delta t}{\Delta t+\tau_{\text{e}}}\boldsymbol{\mathcal{W}}_{\chi}(\boldsymbol{g},\boldsymbol{h}_{\text{rev(p)}})\cdot\tilde{\boldsymbol{\mathcal{W}}}_{\kappa}(\boldsymbol{h},\boldsymbol{h}_{\text{rev(p)}},\boldsymbol{g}_{\text{(p)}}),$ (54)
 where $\boldsymbol{\mathcal{W}}_{\chi}$ is the second-order tensor defined by